### v1

# Towards Better Benchmark Datasets for Inductive Knowledge Graph Completion

Abstract
Knowledge Graph Completion (KGC) attempts to predict missing facts in a Knowledge Graph (KG). Recently, there’s been an increased focus on designing KGC methods that can excel in the inductive setting, where a portion or all of the entities and relations seen in inference are unobserved during training. Numerous benchmark datasets have been proposed for inductive KGC, all of which are subsets of existing KGs used for transductive KGC. However, we find that the current procedure for constructing inductive KGC datasets inadvertently creates a shortcut that can be exploited even while disregarding the relational information. Specifically, we observe that the Personalized PageRank (PPR) score can achieve strong or near SOTA performance on most inductive datasets. In this paper, we study the root cause of this problem. Using these insights, we propose an alternative strategy for constructing inductive KGC datasets that helps mitigate the PPR shortcut. We then benchmark multiple popular methods using the newly constructed datasets and analyze their performance. The new benchmark datasets help promote a better understanding of the capabilities and challenges of inductive KGC by removing any shortcuts that obfuscate performance.
1 Introduction
Knowledge Graph Completion (KGC) attempts to predict unseen facts on an existing knowledge graph (KG). KGC has numerous applications including drug discovery [1], personalized medicine [2], and recommendations [3]. Traditionally, most research on KGC has focused on the transductive setting, where the same entities and relations are seen during training and testing. Most methods [4, 5, 6] generally focus on learning embeddings for all entities and relations to facilitate the prediction of new facts.
In recent years, interest in KGC has shifted towards designing methods that can generalize to new entities or relations not seen during training. This task, known as “inductive KGC”, requires a method to train on a graph $\mathcal{G}_{\text{train}}$ and perform inference on a different graph $\mathcal{G}_{\text{inf}}$, where the inference graph contains either new entities, relations, or both. Because of this, methods for inductive KGC don’t rely on fixed embeddings for entities or relations, instead opting for more flexible techniques that can inductively learn representations based on a given graph [7, 8, 9]. To asses the ability of methods for this task, new datasets have been constructed that require the ability to reason inductively. All inductive datasets [7, 10, 9] are constructed from existing transductive KGC datasets. This is done by sampling two graphs that contain disjoint entities, one each for train and inference. Multiple methods [8, 11, 9] have reported tremendous promise on these newer benchmark datasets.
However, we find that on almost all inductive datasets, we can achieve competitive performance by using the Personalized PageRank [12] (PPR) score to perform inference. We note that PPR is a non-learnable heuristic and ignores the relational information in the graph. In Figure 1, we compare the performance of PPR against the supervised SOTA performance on both inductive and transductive datasets. We can see that when performing KGC on inference graphs with either new entities (E) or new entities and relations (E, R), PPR performs only roughly 25% worse than SOTA. However, this is generally not true for transductive datasets, where PPR usually performs poorly. Interestingly, this observation is true for inductive datasets even when the transductive dataset that it is created from has poor PPR performance. For example, while PPR performs very poorly on FB15k-237, it achieves higher performance on it’s inductive derivatives (denoted by “FB”). These findings are problematic as PPR has no basis in literature as a heuristic for KGC, since it completely overlooks the relational aspect of KGs. Therefore, this suggests the potential existence of a shortcut that allows a simple non-learnable method like PPR to achieve high performance on almost all inductive datasets . This also brings into question how successful most methods are in inductive reasoning, as a large portion of their performance may be due to this shortcut.
This finding naturally motivates us to ask – why can PPR perform so well on existing inductive datasets? In Section 3 we discover that the high performance of PPR is due to how the inductive datasets are created from transductive datasets. Specifically, we observe that the current procedure creates graphs where the shortest path distance (SPD) between entities in positive test samples is much lower than the SPD between those in negative samples. This allows for a method like PPR, which gives a higher weight to shorter walks, to distinguish between the positive and negative samples based solely on the distance. To account for this problem, in Section 4 we propose a new strategy for sampling inductive datasets that uses graph partitioning to create the train and inference graphs. This allows us to sample subgraphs that retain the general properties of the original graph. We then demonstrate that this procedure can indeed create inductive datasets that limit the performance ability of PPR. Lastly, we benchmark common inductive KGC methods on our newly constructed datasets. Our key contributions can be summarized as follows:
  • We observe that on existing inductive KGC datasets, we can achieve competitive performance when ranking entities using only the Personalized PageRank score.
  • Through empirical study, we find that the strong performance of PPR on inductive datasets is due to the current procedure for constructing those datasets, which allows for a shortcut to distinguish between positive and negative samples.
  • We propose a new strategy for sampling inductive KGC datasets from their transductive counterparts that uses graph partitioning. We show that our proposed method can substantially mitigate the shortcut. We then benchmark popular inductive KGC method on our newly created datasets. Compared to the older datasets, previous SOTA tend to decrease in performance.
The remaining paper is structured as follows. In Section 2, we provide background on inductive KGC methods, datasets, and PPR. In Section 3, we study in detail when and why PPR can perform well on inductive KGC. We then introduce our new strategy for creating inductive datasets in Section 4 and benchmark popular methods on these newly created datasets in Section 5.
(a) (E) Inductive
(b) (E, R) Inductive
(c) Transductive
Figure 1: Hits@10 of PPR vs. Supervised SOTA. Results are on (a) (E) Inductive, (b) (E, R) Inductive, and (c) Transductive datasets.
2 Background
Throughout this study we denote a knowledge graph as $\mathcal{G}=\{\mathcal{V},\mathcal{R},\mathcal{E}\}$ where $\mathcal{V}$ is a set of entities (i.e., nodes), $\mathcal{R}$ the set of relations (i.e., edge types), and $\mathcal{E}$ the set of edges (i.e., triples) of the form $(s,r,o)$ where $s$ and $o$ are entities and $r$ a relation. Lastly, we note that the task of KGC is formulated as the following where given a query $(s,r,*)$, we attempt to predict the correct entity *.
Inductive KGC Datasets: In inductive KGC, we are given a training graph $\mathcal{G}_{\text{train}}=\{\mathcal{V}_{\text{train}},\mathcal{R}_{\text{train}},\mathcal{E}_{\text{train}}\}$ and an inference graph $\mathcal{G}_{\text{inf}}=\{\mathcal{V}_{\text{inf}},\mathcal{R}_{\text{inf}},\mathcal{E}_{\text{inf}}\}$. A method is trained on $\mathcal{G}_{\text{train}}$ and evaluated on $\mathcal{G}_{\text{inf}}$. Most datasets consider the setting that is only disjoint on the entities such that $\mathcal{R}_{\text{inf}}\in\mathcal{R}_{\text{train}}$ and $\mathcal{E}_{\text{inf}}\cap\mathcal{E}_{\text{train}}=\emptyset$. This setting is referred to as the (E) setting. Another setting, which we denote as (E, R) further allows $\mathcal{R}_{\text{inf}}$ to contain relations not in $\mathcal{R}_{\text{train}}$.
All existing inductive datasets are sampled from existing transductive datasets. Furthermore, the majority of inductive datasets [7, 9] are further created via the same procedure introduced by [7]. We give brief overview of this procedure below. Given a transductive dataset, which we denote as $\mathcal{G}$, $k$ seed entities are randomly chosen from $\mathcal{G}$. The 2-hop neighborhood is then extracted around each individual seed entity. To prevent exponential growth, the number of neighbors sampled at any hop is capped at 50 for each individual seed entity. The resulting edges are then combined to create $\mathcal{G}_{\text{train}}$ and are subsequently removed from the original graph. The inference graph is then sampled in a similar manner from the resulting graph $\mathcal{G}\setminus\mathcal{G}_{\text{train}}$. We note that the one exception to this procedure are the ILPC datasets, introduced by [10], which instead sample $p\%$ of the nodes from $\mathcal{G}$ and use them to create $\mathcal{G}_{\text{train}}$. The rest of the nodes are then used to construct $\mathcal{G}_{\text{inf}}$. In practice, we find that both sampling methods tend to produce similar graphs.
Inductive KGC Methods: Both NeurlLP [13] and DRUM [14] consider combining the path representations of different length between both entities in a triple. However, since they explicitly consider each path, they are often limited to only considering paths of up to length 2 or 3. Conditional MPNNs [15] are a more efficient alternative to encoding higher-order path information. They work by conditioning the message passing mechanism on the known entity, allowing the implicit aggregation of all paths up to a length $L$ (which is equal to the number of GNN layers). The value of $L$ is typically $5/6$. Prominent examples include NBFNet [8] and RED-GNN [11]. More scalable alternatives have been proposed that prune the propagated messages [16, 17, 18]. NodePiece [19] is concerned with parameter efficiency, using an anchor-based approach to learn a more compact set of entity and relation representations. All previous methods assume that a fixed set of relations exist between the train and inference graphs. To account for new relations in inference, InGRAM [9] introduces the concept of a “relation graph”, which inductively encodes the representation of each relation. Note that we omit subgraph methods that have been to shown to prohibitively expensive [7, 20, 21, 22, 23] or those that require the use of external textual information [24].
Personalized PageRank: PageRank [12] computes the probability of finishing a random walk of arbitrary length at some node $u$, when there is equal probability of beginning the walk at any node in the graph. Personalized PageRank (PPR) [12] is a version of PageRank that is “personalized" to some root node $s$, where at each step there is a probability $\alpha$ of teleporting to $s$. The set of PPR scores for a root node $s$ is given by $\text{pr}_{s}\in\mathbb{R}^{\lvert\mathcal{V}\rvert}$ and can be formulated as the weighted sum of all random walk probabilities between two nodes [25]:
$\text{pr}_{s}=\alpha\sum_{k=0}^{\infty}(1-\alpha)^{k}W^{k}x_{s},$ (1)
where $W$ is the random walk matrix and $x_{s}$ is a one-hot vector at node $s$. Intuitively, we can see that weight given to a walk decays with the increase in length due to $(1-\alpha)^{k}$. As such, the PPR score will often be higher for those nodes of shorter distance to $s$. Note that for a KG, we obtain the PPR matrix by first converting the inference graph, $\mathcal{G}_{\text{inf}}$, to an undirected graph. We note that this is common practice in KGC [26] whereby inverse relations are added to the graph. We further assume an edge weight of 1 for all edges.
3 Preliminary Study
In this section, we examine the performance on common inductive KGC datasets. We first show that the Personalized PageRank (PPR) can often serve as a good baseline on most datasets. From this observation, we attempt to answer two important questions: (1) When can the PPR scores perform well on KGC? and (2) Why can PPR sometimes perform well for KGC?
3.1 Personalized Pagerank (PPR) is a Strong Baseline for Inductive KGC
We begin by obtaining the PPR matrix of the inference graph for use in evaluation. See Section 2 for more details on how this is done. Given this new graph, for a query $(s,r,*)$, we calculate the PPR score for all possible entities $o\in\mathcal{V}$. Using these scores, we can obtain the rank of the true entity for the query and compute the performance in terms of Hits@10.
In Figure 1, we show the performance when using the PPR versus the SOTA performance among supervised methods. We split the datasets by transductive, (E), and (E, R) inductive. We include those datasets most often used in each task, comprising 25 in total. See Appendix B.1 for more details on the datasets chosen. The full set of results can be found in Appendix A.1. We observe that on both types of inductive datasets, the PPR score does reasonably well, performing on average only 25-29% less than SOTA. Also, on some datasets like the WN or ILPC inductive splits, the performance nearly matches or exceeds the supervised performance. On the other hand, for the transductive datasets, the performance disparity is often much larger. Interestingly, PPR still performs well on some transductive datasets, including the popular WN18RR. This tells us that this phenomenon is not unique to inductive datasets. However, this phenomenon is far less common on transductive datasets.
Curiously, for the inductive datasets, their performance using PPR is much higher than their transductive parents. For example, on the inductive datasets derived from FB15k-237, the performance of PPR increases by an average of 1,186%. Furthermore, for those derived from WN18RR, they see an average increase of 43%. This suggests that there is some change in the underlying inductive graphs that are causing the performance of PPR to increase.
3.2 When Does PPR Perform Well and Why?
In the previous section we detailed that PPR score performs well on the inductive setting. Furthermore, the performance of PPR on inductive datasets is much higher than their transductive counterparts of which they’re derived from. This raises the question – why can PPR perform well on some datasets but not others?
Figure 2: $\Delta\text{SPD}$ vs. PPR Hits@10.
We find that the answer lies in comparing the mean shortest path distance (SPD) between positive and negative samples. For a query $(s,r,*)$, let $o^{+}$ denote the correct answer and $o^{-}\in\mathcal{V}^{-}(s,r)$ as the set of negative answers for the query. We compute the SPD on the inference graph for both $o^{+}$ and $\mathcal{V}^{-}(s,r)$. This is repeated for each test sample $(s,q,*)\in\mathcal{E}_{\text{inf}}$. We then calculate the mean SPD across all positive and negative test samples, which we denote as $\text{SPD}^{+}$ and $\text{SPD}^{-}$, respectively. The different in mean SPD is correspondingly given by $\Delta\text{SPD}=\text{SPD}^{+}-\text{SPD}^{-}$. This tells us, on average, how much shorter the distance between entities in positive samples are relative to those in negatives.
In Figure 2 we show the relationship between $\Delta\text{SPD}$ and the Hits@10 when using PPR. The results are across 25 different datasets, for both inductive and transductive KGC. Calculating the Pearson correlation, we can observe it is 0.87, which indicates a strong relationship between the two metrics. Therefore, there exists a basic pattern to distinguish positive and negative samples in many KGC datasets. Put simply, the SPD between entities in positive test samples tend to be lower than that in negative samples. The larger the discrepancy between the mean distances, the better PPR can perform. This suggests that the PPR scores can exploit this distance disparity in the datasets to achieve good performance, even while completely ignoring the relational information.
But, why can PPR exploit this pattern? As shown in Eq. (1), the PPR score between two nodes is the weighted sum of random walk probabilities between them. Furthermore, walks of shorter length are weighed more heavily than those of longer length. Therefore, nodes with a shorter distance between them will be able to benefit from these highly-weighted walks, while those of larger distance will not. For example, when $\alpha=0.15$, the weight of a walk with $\text{SPD}=2$ is $0.72$ while when $\text{SPD}=5$ it is $0.44$. As such, the PPR score will invariably favor those node pairs with a lower SPD.
Note that we use PPR as opposed to SPD in our experiments since all-pairs SPD is costly to calculate and the full PPR matrix can be efficiently approximated via [27]. Furthermore, as we show in Section 3.3, even when controlling for the SPD, the PPR score can help differentiate between positive and negative samples.
3.3 Why is This So Common on Inductive Datasets?
(a) FB15k-237
(b) WN18RR
Figure 3: The mean shortest path (SP) distance between positive test samples and negative samples. We display the results for two transductive datasets (FB15k-237 and WN18RR) and the inductive datasets that are derived from them.
In the last section, we covered when PPR can perform well on KGC and why. However, one remaining question is – why is this trend so pervasive on inductive datasets but rarely on their transductive counterparts? For example, on FB15k-237 [28], a transductive dataset, the Hits@10 of PPR is 2.7%. However, across eight different inductive versions of FB15k-237, the mean PPR performance is 32%. We find that this can be explained by the following two observations.
Observation 1: The current procedure for creating inductive datasets increases the $\Delta\text{SPD}$ and thereby the performance of PPR. As noted earlier, all common inductive datasets are created from existing transductive datasets. See Section 2 for a detailed overview of the construction process. In a nutshell, the training and inference graph are constructed sequentially by sampling a number of subgraphs from a graph. In Figure 3 we show the mean SPD of both positive and negative samples for 12 inductive datasets and their parent transductive dataset. We limit our analysis to those datasets derived from FB15k-237 and WN18RR, as the majority of inductive datasets are derived from them. We observe that for all inductive datasets, while the mean SPD for negative samples sharply rises, the mean SPD for positive samples stay roughly the same. This creates the shortcut described in Section 3.2, where the SPD can easily differentiate between positive and negative samples. Since the gap between the distances is almost always larger than the original transductive dataset, this shortcut becomes more pronounced, thereby leading to a better PPR performance (see Appendix A.1 for detailed results).
But why does the mean SPD drastically change for negative samples but not positive? We find that this is due to the fact that entities in positive samples are more well-connected those those in negatives. In Table 1, we show the % difference in mean PPR score between positive and negative samples when controlling for the SPD on two inductive datasets.
Table 1: % Increase in in Pos over Neg PPR, controlling for SPD.
SPD Range WN18RR v4 FB15k-237 v4
$[1,2)$ +17% +5%
$[2,3)$ +29% +200%
$[3,4)$ +82% +328%
$[4,\infty)$ +2837% +44%
We can see that the PPR score is much greater for positive samples, even for higher values of SPD. Therefore, the SPD of positive samples are better able to “withstand" changes in the underlying graph better than negatives, as they typically contain additional shorter walks between samples. However, negatives samples are typically much less well-connected, so they are more affected by removing a portion of edges from the original graph.
(a) $\Delta\text{SPD}$ vs. PPR Hit@10
(b) PPR Hit@10 vs. Train Size
(c) PPR Hit@10 vs. Inference Size
Figure 4: Experiments when generating inductive datasets from FB15k-237. (a) The generated datasets always have a much higher $\Delta\text{SPD}$ than the transductive dataset. (b)/(c) There is a trade-off between the size of the train and inference graph, making it difficult to obtain a good split.
Observation 2: Constructing a good inductive dataset is difficult. In Section 2 we discuss the general algorithm used to create inductive datasets. Multiple parameters exist that guide the construction process. It is tempting to think that by simply trying different combinations, one can happen upon a split that doesn’t suffer from high PPR performance. However, in practice, we find that this is difficult, as there are multiple factors to contend with.
We demonstrate this by attempting to generate inductive datasets from FB15k-237. We generate a number of different inductive datasets by modifying the (a) # of seed entities for train and inference and (b) the maximum neighborhood size for train and inference. For different combination of values, we generate three different datasets using different random seeds. For each of the generated datasets, we calculate the size of both the train and inference graphs, the $\Delta\text{SPD}$, and the PPR performance. These values are then averaged across seeds. A more detailed discussion is given in Appendix C. In Figure 4(a) we plot the $\Delta\text{SPD}$ vs. the PPR performance. Despite searching across a wide variety of parameters, both $\Delta\text{SPD}$ and the PPR performance remain much higher for the inductive datasets compared to the transductive dataset. Furthermore, we show the relationship between the size of the train and inference graphs and the PPR performance in Figures 4(b) and 4(c). We observe that when the performance of PPR is at its lowest, the size of the train graph is noticeably small. However, fixing this problem, results in a sharp increase in the performance of PPR. Furthermore, there is an inverse relationship between the size of the train and inference graph, making it hard to find a “sweet spot” where the PPR performance is low and both graphs aren’t too small.
The studies in this section show that the current strategy for constructing inductive datasets from existing transductive datasets is liable to introduce a well-performing shortcut into the existing graph. Currently, it almost always leads to a sharp increase in the PPR performance. Furthermore, attempting to limit the severity of this issue is very difficult while also generating train and inference graphs of reasonable sizes. This suggests that a new strategy is needed for sampling graphs for inductive KGC from transductive datasets.
4 Constructing Inductive Datasets through Graph Partitioning
In the previous section, we showed that the PPR score can achieve strong performance on most inductive KGC datasets. Furthermore, we demonstrate that this is due to how inductive datasets are sampled from existing transductive datasets. This sampling strategy engenders a shift in the underlying properties of the graph that allows for PPR to perform well. This naturally causes us to ask – How can we mitigate this problem when constructing newer inductive datasets? In the next subsection, we introduce our strategy which utilizes graph partitioning to alleviate this problem.
4.1 Partition-Based Dataset Sampling
We’ve previously covered in Section 3 that the existing procedure for creating inductive datasets leads to suboptimal subgraphs. This is because the resulting subgraphs tend to have much different properties than the original graph, such as the distance distribution, which can potentially lead to shortcuts when performing KGC.
We note that the task of constructing inductive datasets from an existing graph $\mathcal{G}$ can be framed as a graph partitioning problem. Formally, we want to sample two non-overlapping partitions from the graph such that $\mathcal{G}_{\text{train}},\mathcal{G}_{\text{inf}}\in\mathcal{G}$. Given the analysis in Section 3.2, we hope to sample subgraphs such that $\Delta\text{SPD}(\mathcal{G}_{\text{train}})\approx\Delta\text{SPD}(\mathcal{G}_{\text{inf}})\approx\Delta\text{SPD}(\mathcal{G}$). But, how do we find subgraphs that satisfy this property? Intuitively, we want to sample each subgraph so that it’s removal has little effect on the initial graph’s structure. We give an example in Figure 5 where we sample two subgraphs from an existing graph. As we can see, even though the graph is partitioned, the relationship between entities in the same partition remain roughly the same before and after the partition. This is because there already exists little relationship between the two partitions in the original graph.
Table 2: $\Delta\text{SPD}$ and PPR Performance of New and Old Inductive Datasets vs. Transductive Parents.
Task Dataset PPR Hits@10 $\Delta\text{SPD}$
Trans. Ind. New Ind. Old Trans. Ind. New Ind. Old
(E) WN18RR 46.2 45.1 66.0 4.1 3.2 6.3
CoDEx-M 9.0 11.2 21.1 0.2 0.24 0.78
HetioNet 2.7 2.4 NA 0.55 0.29 NA
(E, R) FB15k-237 2.7 10.8 21.4 0.46 0.48 2.42
CoDEx-M 9.0 13.2 NA 0.20 0.42 NA
Figure 5: Sampling Two Subgraphs
Multiple popular approaches [29, 30] exist that attempt to divide the graph into optimal partitions. The guiding principle in these approaches is that the partitions should be internally dense but only sparsely connected to one another. Because of this, the entities in different partitions should only be weakly connected and have little impact on one another. Therefore, the relationship between entities in the same partition are minimally affected by outside entities or edges. As such, removing this partition from the graph should then have little effect on the entities in that partition. Also, since the partitions are created at the same time, we avoid sampling the inference graph after the training, which can be suboptimal.
In practice, we consider using Spectral Clustering [29] or the Louvain method [30], as dependent on the dataset. Once the original graph is partitioned into $n$ partitions, we sample $k$ of those to be used. The partitions are chosen such that they display similar properties to the original graph (see Section 4.2 for more). Of the $k$ partitions, one is chosen as the training graph while the other $k-1$ are designated as separate inference graphs. This is an important advantage of using partitioning to sample the graphs, as it allows us to avoid having to sample multiple different train and inference pairs as in [7, 9], thereby allowing for more efficient benchmarking. Lastly, we note that when sampling graphs for the (E) inductive task, we further remove new relations from the inference graph. In practice, we find that we can easily find partitions where this amounts to little or no change in the inference graph. See Section B.2 for more details on the dataset creation process.
4.2 Analysis of New Datasets
In this section we analyze the new inductive datasets created following the partition-based procedure outlined in Section 4.1. We create (E) inductive datasets, from WN18RR [26], CoDEx-M [31], and HetioNet [32]. For the (E, R) datasets, we use CoDEx-M [31] and FB15k-237 [28]. Each dataset has two separate inference graph except for CoDEx-M on the (E) task which has one. Note that some datasets are only suitable for one task or another. For example, WN18RR and HetioNet have very few relations, making it nearly impossible to sample two partitions with significantly different relations for the (E, R) task. On the other hand, FB15k-237 contains too many relations to sample multiple graphs for the (E) task without removing many edges from either the train or inference graph.
In Table 2 we show the PPR Hits@10 and $\Delta\text{SPD}$ for the inductive datasets and their original transductive dataset. We note that when multiple inference graphs exist, we take the mean across each inference graph. When possible, we also include a comparison against those inductive datasets that already exist. For example, for WN18RR in the (E) task, 4 datasets exist from [7]. As compared to the old inductive datasets, the PPR performance for the new inductive datasets is much lower. On average, the performance of PPR is 78% lower for the new inductive datasets. Furthermore, the PPR performance of the new inductive datasets is very similar to the performance on the original transductive dataset. A similar trend can be found when comparing the $\Delta\text{SPD}$. This analysis shows that newer sampling procedure can indeed sample inductive datasets that are much more similar to the original transductive graph. Furthermore, it greatly helps mitigate the PPR shortcut.
5 Experiments
5.1 Experimental Settings
Datasets: We use the new datasets created in Section 4.1. For the (E) setting, this includes WN18RR, HetioNet, and CoDEx-M. For the (E, R) setting, it is FB15k-237 and CoDEx-M. We sample 2 inference graphs for each dataset, except for CoDEx-M on the (E) setting where we could only find 1 suitable graph for inference. For each inference graph, 10% of edges are randomly removed for testing. For validation 10% of edges are removed from the training graph. It is necessary that the validation samples are extracted from the train graph as the inference graphs must remain unobserved during training. The statistics for each dataset can be found in Appendix B.2.
Baseline Methods: We consider prominent KGC methods including NBFNet [8], RED-GNN [11], NodePiece [19], and InGram [9]. We also consider the recent foundation model ULTRA [33].
Hyperparameters: The exact hyperparameters and their ranges differ by method due to difference in implementations. For each method, we tried to follow the recommended ranges used by the authors. For InGram [9] we tuned the learning rate in $\{1e^{-3},5e^{-4}\}$ and the number of entity layers in $\{2,4\}$. For NodePiece [19] we tuned the learning rate from $\{1e^{-3},1e^{-4}\}$ and the loss margin from $\{15,25\}$. For RED-GNN [11], we tuned the learning rate from $\{5e^{-3},5e^{-4}\}$ and the dropout from $\{0.1,0.3\}$. In their experiments, NBFNet [8] uses the same hyperparameters across all datasets. We therefore do the same.
The full set of experimental settings can be found in Appendix D.
5.2 Results
Table 3: (E) Inductive Results (Hits@10) for supervised methods.
Models CoDEx-M WN18RR HetioNet
Inference 1 Inference 1 Inference 2 Inference 1 Inference 2
PPR 11.2 66.2 24 3.2 2.2
NodePiece 6.8 ± 0.8 29.6 ± 0.8 4.8 ± 0.6 10.2 ± 0.9 15.4 ± 0.9
InGram 20.1 ± 3.5 38.0 ± 2.4 8.0 ± 2.9 21.9 ± 1.1 22.3 ± 2.8
RED-GNN 35.6 ± 2.3 72.9 ± 0.4 27.7 ± 0.3 68.3 ± 3.0 85.1 ± 2.7
NBFNet 43.6 ± 0.2 75.5 ± 0.2 29.4 ± 2.5 72.8 ± 3.8 77.2 ± 0.4
Table 4: (E, R) Inductive Results for supervised methods. The % of new relations are in parentheses.
Models FB15k-237 CoDEx-M Inference 1 (27%) Inference 2 (63%) Inference 1 (10%) Inference 2 (57%) MRR Hits@10 MRR Hits@10 MRR Hits@10 MRR Hits@10 PPR 3.8 9.1 4.7 12.4 3.3 10.9 7.1 15.4 NodePiece 1.4 ± 0.1 3.0 ± 0.6 2.6 ± 0.2 4.7 ± 0.5 1.5 ± 0.2 3.1 ± 0.6 1.6 ± 0.4 2.5 ± 1.0 InGram 11.2 ± 1.9 23.8 ± 3.0 9.2 ± 0.9 20.2 ± 2.0 10.3 ± 1.2 20.4 ± 3.3 7.2 ± 4.2 15.9 ± 10.0 RED-GNN 11.8 ± 2.8 21.6 ± 5.8 19.8 ± 2.2 33.3 ± 4.2 16.8 ± 1.6 29.2 ± 2.9 14.7 ± 5.3 26.5 ± 10.4 NBFNet 15.8 ± 0.9 27.5 ± 1.8 15.9 ± 0.2 26.2 ± 0.3 27.9 ± 8.8 47.7 ± 11.8 8.5 ± 4.8 17.6 ± 10.0
Main Results: The main results for supervised methods can be found in Tables 3 and 4. We observe that on nearly every dataset, NBFNet and RED-GNN are the two best models. This indicates that conditional MPNNs [15], the class of model in which both belong to, are necessary for strong performance in inductive KGC. Interestingly, we observe that InGram struggles in the (E, R) setting. This is even true when the % of new relations is high. This run counters to the results on older inductive datasets [9] where InGram excelled over NBFNet and RED-GNN. Lastly, outside of NodePiece, each method is usually able to significantly outperform PPR.
Performance Comparison on New and Old Inductive Datasets: We compare the change in performance from the previous inductive datasets to our new datasets. Note that we only compare when there is a corresponding dataset sampled from the same transductive dataset and for the same task. For the (E) task, this includes CoDEx-M (ILPC-S [10]) and WN18RR (v1-v4 [7]). For the (E, R) task it is only FB15k-237 (25-100 [9]). For each, we compare the performance of the SOTA on the older dataset to their corresponding performance on the new datasets. This is NBFNet [8] for WN18RR and FB15k-237 and NodePiece [19] for CoDEx-M. We only compare the performance of the old SOTA as not all methods were tested on the older datasets or guranteed to be well-tuned. The results are shown in Figure 6(a). We find that performance on each of the new datasets drops significantly. This suggests that removing the shortcut has a large negative effect on the performance.
Performance of ULTRA [33]: We further compare against ULTRA, a recent foundation model designed for fully inductive KGC. We evaluate ULTRA under the 0-shot setting. See Appendix E for more details on the version of ULTRA used. The results are in Figure 6(b) where for each dataset, we average the results across the different inference graphs (full results in Appendix A.2). On the (E) task, ULTRA is comparable to the supervised SOTA. However, on the (E, R) task, ULTRA significantly outperforms other methods. This suggests that ULTRA contains a much greater ability to generalize to unseen relations than existing supervised methods.
(a) Performance of SOTA on Old vs. New Datasets
(b) Results across Inf. graphs for ULTRA vs. SOTA
Figure 6: (a) Performance of SOTA on old vs. new datasets, (b) ULTRA [33] vs. supervised SOTA.
6 Conclusion
In this paper we study the problem of constructing datasets for inductive knowledge graph completion. Upon examination, we find that we can achieve competitive performance on most inductive datasets through the use of Personalized PageRank [12], which ignores the relational structure of the graph. Through our study, we uncover that this shortcut is due to how inductive datasets are created. To remedy this problem, we propose a new dataset construction process based on graph partitioning that helps mitigate the impact of the studied shortcut. We then construct new benchmark datasets using this new procedure and benchmark various methods.
Appendix A Additional Results
A.1 Supervised SOTA vs. PPR Performance
In this section we show the performance in terms of Hits@10 for the SOTA supervised method vs. the Personalized PageRank (PPR) score. We further include the % difference in performance and which method is considered SOTA. These are shown in Tables 56, and 7 for the (E), (E, R), and transductive datasets, respectively. For an overview of SOTA performance on additional KG datasets, please see [33].
Table 5: Hits@10 for PPR vs. Supervised SOTA on the (E) Inductive Datasets
Dataset Supervised SOTA PPR % Difference SOTA Method
WN v1 82.6 77.1 7% NBFNet [8]
WN v2 79.8 74.4 12% NBFNet [8]
WN v3 56.8 45.2 26% NBFNet [8]
WN v4 74.3 67.3 10% A*Net [16]
FB v1 60.7 41.2 47% NBFNet [8]
FB v2 70.4 47.6 48% NBFNet [18]
FB v3 66.7 43.5 53% NBFNet [8]
FB v4 66.8 38.4 74% NBFNet [8]
ILPC-S 25.1 19.8 27% NodePiece [19]
ILPC-L 14.6 22.5 -35% NodePiece [19]
Table 6: Hits@10 for PPR vs. Supervised SOTA on the (E, R) Inductive Datasets
Dataset Supervised SOTA PPR % Difference SOTA Method
FB-100 37.1 22.2 67% InGram [9]
FB-75 32.5 21.9 48% InGram [9]
FB-50 21.8 20.5 6% InGram [9]
FB-25 27.1 20.9 30% InGram [9]
WK-100 16.9 15.8 7% InGram [9]
WK-75 36.2 29.5 23% InGram [9]
WK-50 13.5 10.6 27% InGram [9]
WK-25 30.9 23.2 33% InGram [9]
Table 7: Hits@10 for PPR vs. Supervised SOTA on the Transductive Datasets
Dataset Supervised SOTA PPR % Difference SOTA Method
FB15k-237 66.6 2.7 2367% NBF+TAGNet [18]
WN18RR 59.9 46.2 30% NBF+TAGNet [18]
CoDEx-M 49.0 9.0 444% ComplEx RP [34]
CoDEx-S 66.3 8.6 671% ComplEx RP [34]
CoDEx-L 47.3 9.0 426% ComplEx RP [34]
Hetionet 40.3 2.4 1579% RotatE [35]
DBPedia100k 41.8 30.2 38% ComplEx-NNE+AER [36]
A.2 Performance of ULTRA
We include the full results of ULTRA [33] on each dataset and inference graph. The result on the (E) datasets are in Table 8 while those for the (E, R) datasets are in Table 9.
Table 8: (E) Inductive Results for ULTRA [33]
Metric CoDEx-M WN18RR HetioNet
Inference 1 Inference 1 Inference 2 Inference 1 Inference 2
MRR 30.2 64.7 21.4 57.9 72.7
Hits@10 46.6 72.7 38.5 69.1 86.3
Table 9: (E) Inductive Results for ULTRA [33]
Metric FB15k-237 CoDEx-M
Inference 1 Inference 2 Inference 1 Inference 2
MRR 45.7 38 30.4 73.7
Hits@10 69.6 64.5 45.7 91.3
Appendix B Datasets
B.1 Existing Datasets
We detail the statistics of all existing transductive and inductive datasets in Tables 10, 11, and 12, respectively. We further include the licenses for each in Table 13. Note that we omit YAGO3-10 [37] as [38] show that the dataset is dominated by two duplicate relations, making most triples trivial to classify. Also, we omit NELL-995 and any inductive datasets derived from it due to findings by [31] that show that most triples in the dataset are either too generic or meaningless.
Table 10: Statistics for Transductive Datasets.
Dataset #Entities #Relations #Train #Validation #Test
FB15k-237 [28] 14,541 237 272,115 17,535 20,466
WN18RR [26] 40,943 11 86,835 3,034 3,134
CoDEx-S [31] 2,034 42 32,888 1,827 1,828
CoDEx-M [31] 17,050 51 185,584 10,310 10,311
CoDEx-L [31] 77,951 69 551,193 30,622 30,622
HetioNet [32] 45,158 24 2,025,177 112,510 112,510
DBpedia100k [36] 99,604 470 597,572 50,000 50,000
Table 11: Statistics for (E) Inductive Datasets.
Dataset #Relations Train Graph Validation Graph Test Graph
#Entities #Triples #Entities #Triples #Valid #Entities #Triples #Test
FB15k-237 v1 [7] 180 1,594 4,245 1,594 4,245 489 1,093 1,993 205
FB15k-237 v2 [7] 200 2,608 9,739 2,608 9,739 1,166 1,660 4,145 478
FB15k-237 v3 [7] 215 3,668 17,986 3,668 17,986 2,194 2,501 7,406 865
FB15k-237 v4 [7] 219 4,707 27,203 4,707 27,203 3,352 3,051 11,714 1,424
WN18RR v1 [7] 9 2,746 5,410 2,746 5,410 630 922 1,618 188
WN18RR v2 [7] 10 6,954 15,262 6,954 15,262 1,838 2,757 4,011 441
WN18RR v3 [7] 11 12,078 25,901 12,078 25,901 3,097 5,084 6,327 605
WN18RR v4 [7] 9 3,861 7,940 3,861 7,940 934 7,084 12,334 1,429
ILPC-S [10] 48 10,230 78,616 6,653 20,960 2,908 6,653 20,960 2902
ILPC-L [10] 65 46,626 202,446 29,246 77,044 10,179 29,246 77,044 10,184
Table 12: Statistics for (E, R) Inductive Datasets.
Dataset Train Graph Validation Graph Test Graph
#Entities #Rels #Triples #Entities #Rels #Triples #Valid #Entities #Rels #Triples #Test
FB-25 [9] 5,190 163 91,571 4,097 216 17,147 5,716 4,097 216 17,147 5,716
FB-50 [9] 5,190 153 85,375 4,445 205 11,636 3,879 4,445 205 11,636 3,879
FB-75 [9] 4,659 134 62,809 2,792 186 9,316 3,106 2,792 186 9,316 3,106
FB-100 [9] 4,659 134 62,809 2,624 77 6,987 2,329 2,624 77 6,987 2,329
WK-25 [9] 12,659 47 41,873 3,228 74 3,391 1,130 3,228 74 3,391 1,131
WK-50 [9] 12,022 72 82,481 9,328 93 9,672 3,224 9,328 93 9,672 3,225
WK-75 [9] 6,853 52 28,741 2,722 65 3,430 1,143 2,722 65 3,430 1,144
WK-100 [9] 9,784 67 49,875 12,136 37 13,487 4,496 12,136 37 13,487 4,496
Table 13: Dataset Licenses.
Datasets License
FB15k-237 CC-BY-4.0
WN18RR MIT License
CoDEx-S/M/L MIT License
HetioNet CC0 1.0 Universal
DBpedia100k None
FB15k-237 v1/v2/v3/v4 None
WN18RR v1/v2/v3/v4 None
ILPC-S/L MIT License
FB-25/50/75/100 None
WK-25/50/75/100 None
B.2 New Datasets
The raw data for the new datasets introduced in Section 4, can be found at https://github.com/HarryShomer/Better-Inductive-KGC/tree/master/new_data. Please see the README for more information on how to load the data.
We further detail the statistics of all the new datasets in Tables 14 and 15.
Table 14: Statistics for New (E) Inductive Datasets.
Dataset Graph # Edges # Entities # Rels # Valid/Test $\Delta\text{SPD}$
CoDEx-M Train 76,960 8,362 47 8,552 NA
Inference 1 69,073 8,003 40 7,674 0.24
WN18RR Train 24,584 12,142 11 2,458 NA
Inference 1 18,258 8,660 10 1,831 4.87
Inference 2 5,838 2975 8 572 1.47
HetioNet Train 101,667 3,971 14 11,271 NA
Inference 1 49,590 2,279 11 5,490 0.38
Inference 2 37,927 2455 12 4,187 0.19
Table 15: Statistics for New (E, R) Inductive Datasets.
Dataset Graph # Edges # Entities # Rels # Valid/Test % New Rels $\Delta\text{SPD}$
FB15k-237 Train 45,597 2,869 105 5,062 NA NA
Inference 1 35,937 1,835 143 3,992 62.8% 0.63
Inference 2 51,693 2,606 143 5,735 27.1% 0.34
CoDex-M Train 29,634 4,038 36 3293 NA NA
Inference 1 70,137 7,938 39 7,794 9.9% 0.24
Inference 2 8,821 2,606 28 979 56.8% 0.60
Appendix C Inductive Generation Experiments
We give more details on the experiments conducted in Section 3.3 and shown in Figures 4. Given the transductive dataset FB15k-237 [28], we generate a number of different inductive datasets using the common procedure used by [7, 9]. We describe this procedure in detail in Section 2. Note that for simplicity, we only created datasets for the (E) inductive task. However, the same conclusions should hold for (E, R). Lastly, we only create one graph for inference as is done in previous work.
We generate the inductive datasets by modifying the following set of parameters:
  • # of Initial Train Entities: This is the number of entities that are first assigned to the train graph.
  • # of Initial Inference Entities: This is the number of entities that are first assigned to the inference graph.
  • Max Train Neighborhood Size: We extract the 2-hop neighborhood for each of the initial train entities. To avoid exponential growth, we limit the number of entities selected in each hop to 50. For example, if node 1 has 30 1-hop neighbors and 120 2-hop neighbors, then $30+50=80$ entities are added through the expansion of node 1.
  • Max Inference Neighborhood Size: This follows the same logic as for train but for inference.
The default neighborhood size is set to 50 for the inductive datasets created by [7] and [9]. For [9] they consider 10 and 20, initial entities for train and inference, respectively. This information is not reported for [7].
To simulate the effect of the different parameter values, we create inductive datasets by using a number of different combinations of parameters. For each parameter, we modify it while holding the others constant. This allows us to explore the effect of just that variable. It allows us to avoid running an excessive amount of experiments. The default values are those used in [9] and noted above. In total, there are 17 parameter configurations. For each parameter configuration, we create 3 different datasets through the use of 3 different random seeds. This results in a total of 51 inductive datasets. We then evaluate the performance of PPR on each inference graph and calculate the $\Delta\text{SPD}$. The results for those datasets with the same configuration are then averaged together. All possible configurations and their resulting mean statistics can be found in Table 16. These results are visualized in the main content in Figures 4.
Table 16: FB15k-237 - Inductive Generative Experiments Configurations. Results are over 3 Random Seeds.
# Train Ents # Inf. Ents Max Train Max Inf. # Train Edges # Test Edges $\Delta\text{SPD}$ PPR Hits@10
10 20 10 50 7,599 72,688 1.14 0.166
10 20 15 50 18,518 49,057 1.31 0.180
10 20 25 50 43,077 36,054 1.46 0.213
10 20 50 50 91,843 18,370 1.71 0.309
10 20 100 50 129,782 15,179 1.72 0.235
10 20 50 10 91,843 4,304 2.15 0.400
10 20 50 25 91,843 13,374 1.79 0.319
10 20 50 50 91,843 18,370 1.71 0.309
10 20 50 100 91,843 21,512 1.71 0.310
10 20 50 50 91,843 18,370 1.71 0.309
20 20 50 50 149,635 8,258 2.38 0.317
40 20 50 50 210,109 973 3.89 0.408
10 10 50 50 91,843 11,177 1.6 0.325
10 20 50 50 91,843 18,370 1.71 0.309
10 40 50 50 91,843 36,353 1.81 0.266
10 80 50 50 91,843 47,419 1.83 0.254
10 160 50 50 91,843 55,233 1.93 0.259
Appendix D Additional Experimental Settings
Training Settings: Each model was trained on either a: Tesla V100 32Gb, NVIDIA RTX A6000 48Gb, NVIDIA RTX A5000 24Gb, or Quadro RTX 8000 48Gb. All models were implemented in PyTorch [39] and Torch-Geometric [40].
Evaluation: For a given test triple $(s,r,o)$, we strive to predict both $s$ and $o$ individually. This is framed as a ranking problem where we want the probability of the correct entity (e.g., $o$) to rank higher than all possible negative entities. We can calculate the rank of the true entity over all negatives by the following where $p(\cdot)$ denotes the model probability for a triple:
[…]
We note that a lower rank, indicates better performance. In practice, some of the negative entities will be true triples observed during training. As such, following [4] we use the filtered setting and remove those entities from the ranking. Once we obtain the rank of the true entity (i.e., “how many negative entities does it have a higher probability than”) we compute two metrics – the mean reciprocal rank (MRR) and Hits@K. MRR is given by $1/\text{Rank}$ such that when Eq. 2 holds true for $\forall v\in\mathcal{V}$ it will be equal to 1. Hits@K computes whether $\text{Rank}\leq K$.
Appendix E ULTRA Settings
When evaluating ULTRA [33] we use the 0-shot setting. By default, we use the checkpoint they provided that was trained on the transductive datasets: FB15k-237 [28], WN18RR [26], CoDEx-M [31], and NELL-995 [41]. However, since we create our own inductive splits from FB15k-237, CoDEx-M, and WN18RR, there is risk of test leakage. That is, some triples that may have been in the transductive training graph, may now be a test sample in one of our splits. This gives ULTRA an unfair advantage as it has already seen that triple.
To combat this issue, we train three different versions of ULTRA that omits the transductive dataset used to created that specific inductive dataset. This includes:
  • w/o FB15k-237: Trained using WN18RR, CoDEx-M, and NELL-995.
  • w/o WN18RR: Trained using FB15k-237, CoDEx-M, and NELL-995.
  • w/o CoDEx-M: Trained using FB15k-237, WN18RR, and NELL-995.
We follow the same settings used to train the original checkpoints provided by the authors of ULTRA. Lastly, since HetioNet is not one of the four datasets used, no additional model is needed.
The results are shown in Table 17, where the “W/o” column represents the results when removing the parent transductive dataset and “With” corresponds to the original pre-trained checkpoint provided by the authors of ULTRA [33]. In practice, we observe a small but noteworthy decrease in performance when removing the offending transductive dataset from the pre-trained model.
Table 17: Performance (Hits@10) of ULTRA when pre-trained and w/o the parent transductive dataset. Note that this doesn’t apply to HetioNet.
Dataset Inference Graph W/o With % Difference
CoDEx-M (E) 0 46.6 50.2 -7.2%
WN18RR (E) 0 71.9 72.7 +1.1%
1 38.8 38.5 -0.8%
FB15k-237 (E, R) 0 69.6 72.1 -4.5%
1 64.5 64.8 -0.5%
CoDEx-M (E) 0 45.7 50.4 -9.3%
1 91.1 91.3 -0.2%
Appendix F Implementation
All code used for experiments in this study can be found in the following repository – https://github.com/HarryShomer/Better-Inductive-KGC. Please see the README in the repository for more information on how to run the different experiments.
Appendix G Limitations
One potential limitation of our study is that the new inductive datasets are still being created from existing transductive datasets. While our method can help create better and more realistic inductive scenarios, this is limiting as we are still reliant on good quality transductive datasets. Future work can focus on creating inductive KGC datasets directly from existing large-scale knowledge graphs, thereby bypassing the need to sample from existing transductive datasets.
Appendix H Impact Statement
Our method contributed positively to the field of knowledge graph reasoning. By introducing newer and better inductive datasets, we better align the task of KGC to it’s real-world applications. This is essential, because if we want to perform inductive KGC in real-world tasks, we have to be certain that the methods actually work well. As such, the new datasets can serve as a stronger barometer of actual performance on inductive KGC and will help spur the development of newer and more effective KGC techniques. Since KGC has applications in many different fields including question answering, biology, and recommender systems, there is a strong need for methods that can perform well. We have also carefully considered the broader impact of our work from different perspectives and find that there is no apparent risk.

### v2

@@
 1 Introduction
-Knowledge Graph Completion (KGC) attempts to predict unseen facts on an existing knowledge graph (KG).
+Knowledge Graph Completion (KGC) attempts to predict unseen facts given an existing knowledge graph (KG).
 KGC has numerous applications including drug discovery [1], personalized medicine [2], and recommendations [3].
-Traditionally, most research on KGC has focused on the transductive setting, where the same entities and relations are seen during training and testing.
+Traditionally, most research on KGC was focused on the transductive setting, where the same sets of entities and relations are seen during training and testing.
 Most methods [4, 5, 6] generally focus on learning embeddings for all entities and relations to facilitate the prediction of new facts.
@@
 Because of this, methods for inductive KGC don’t rely on fixed embeddings for entities or relations, instead opting for more flexible techniques that can inductively learn representations based on a given graph [7, 8, 9].
-To asses the ability of methods for this task, new datasets have been constructed that require the ability to reason inductively.
+To asses the ability of methods for this task, new datasets have been constructed that require methods to reason inductively.
 All inductive datasets [7, 10, 9] are constructed from existing transductive KGC datasets.
-This is done by sampling two graphs that contain disjoint entities, one each for train and inference.
+This is done by sampling two graphs, one each for train and inference, which contain disjoint entities.
 Multiple methods [8, 11, 9] have reported tremendous promise on these newer benchmark datasets.
@@
 This allows us to sample subgraphs that retain the general properties of the original graph.
-We then demonstrate that this procedure can indeed create inductive datasets that limit the performance ability of PPR.
+We then demonstrate that this procedure can indeed create inductive datasets that greatly mitigates the ability of PPR.
 Lastly, we benchmark common inductive KGC methods on our newly constructed datasets.
@@
   • We observe that on existing inductive KGC datasets, we can achieve competitive performance when ranking entities using only the Personalized PageRank score.
-  • Through empirical study, we find that the strong performance of PPR on inductive datasets is due to the current procedure for constructing those datasets, which allows for a shortcut to distinguish between positive and negative samples.
+  • Through empirical study, we find that the strong performance of PPR on inductive datasets is due to the current procedure for constructing inductive datasets, which allows for a shortcut to distinguish between positive and negative samples.
   • We propose a new strategy for sampling inductive KGC datasets from their transductive counterparts that uses graph partitioning.
 We show that our proposed method can substantially mitigate the shortcut.
-We then benchmark popular inductive KGC method on our newly created datasets.
-Compared to the older datasets, previous SOTA tend to decrease in performance.
+We then benchmark popular methods on our newly created datasets.
+Compared to the older datasets, methods tend to decrease in performance.
 The remaining paper is structured as follows.
-In Section 2, we provide background on inductive KGC methods, datasets, and PPR.
+In Section 2, we provide background on inductive KGC methods and datasets and PPR.
 In Section 3, we study in detail when and why PPR can perform well on inductive KGC.
@@
 Figure 1: Hits@10 of PPR vs. Supervised SOTA. Results are on (a) (E) Inductive, (b) (E, R) Inductive, and (c) Transductive datasets.
-2 Background
-Throughout this study we denote a knowledge graph as $\mathcal{G}=\{\mathcal{V},\mathcal{R},\mathcal{E}\}$ where $\mathcal{V}$ is a set of entities (i.e., nodes), $\mathcal{R}$ the set of relations (i.e., edge types), and $\mathcal{E}$ the set of edges (i.e., triples) of the form $(s,r,o)$ where $s$ and $o$ are entities and $r$ a relation.
+2 Background and Related Work
+Throughout this study we denote a knowledge graph as $\mathcal{G}=\{\mathcal{V},\mathcal{R},\mathcal{E}\}$ where $\mathcal{V}$ are the set of entities (i.e., nodes), $\mathcal{R}$ the set of relations (i.e., edge types), and $\mathcal{E}$ the set of edges (i.e., triples) of the form $(s,r,o)$ where $s$ and $o$ are entities and $r$ a relation.
 Lastly, we note that the task of KGC is formulated as the following where given a query $(s,r,*)$, we attempt to predict the correct entity *.
@@
 A method is trained on $\mathcal{G}_{\text{train}}$ and evaluated on $\mathcal{G}_{\text{inf}}$.
-Most datasets consider the setting that is only disjoint on the entities such that $\mathcal{R}_{\text{inf}}\in\mathcal{R}_{\text{train}}$ and $\mathcal{E}_{\text{inf}}\cap\mathcal{E}_{\text{train}}=\emptyset$.
+Most datasets consider the setting that is only disjoint on the entities such that $\mathcal{R}_{\text{inf}}\subseteq\mathcal{R}_{\text{train}}$ and $\mathcal{E}_{\text{inf}}\cap\mathcal{E}_{\text{train}}=\emptyset$.
 This setting is referred to as the (E) setting.
@@
 Furthermore, the majority of inductive datasets [7, 9] are further created via the same procedure introduced by [7].
-We give brief overview of this procedure below.
+We now give a brief overview of this procedure.
 Given a transductive dataset, which we denote as $\mathcal{G}$, $k$ seed entities are randomly chosen from $\mathcal{G}$.
 The 2-hop neighborhood is then extracted around each individual seed entity.
-To prevent exponential growth, the number of neighbors sampled at any hop is capped at 50 for each individual seed entity.
+To prevent exponential growth, the number of neighbors sampled at any hop is capped at 50 for each seed entity.
 The resulting edges are then combined to create $\mathcal{G}_{\text{train}}$ and are subsequently removed from the original graph.
 The inference graph is then sampled in a similar manner from the resulting graph $\mathcal{G}\setminus\mathcal{G}_{\text{train}}$.
-We note that the one exception to this procedure are the ILPC datasets, introduced by [10], which instead sample $p\%$ of the nodes from $\mathcal{G}$ and use them to create $\mathcal{G}_{\text{train}}$.
+One exception to this procedure are the ILPC datasets, introduced by [10], which instead sample $p\%$ of the nodes from $\mathcal{G}$ and use them to create $\mathcal{G}_{\text{train}}$.
 The rest of the nodes are then used to construct $\mathcal{G}_{\text{inf}}$.
-In practice, we find that both sampling methods tend to produce similar graphs.
-Inductive KGC Methods: Both NeurlLP [13] and DRUM [14] consider combining the path representations of different length between both entities in a triple.
+In practice, we find that both methods tend to produce similar graphs.
+Inductive KGC Methods: NeurlLP [13] and DRUM [14] consider combining the path representations of different length between both entities in a triple.
 However, since they explicitly consider each path, they are often limited to only considering paths of up to length 2 or 3.
@@
 To account for new relations in inference, InGRAM [9] introduces the concept of a “relation graph”, which inductively encodes the representation of each relation.
-Note that we omit subgraph methods that have been to shown to prohibitively expensive [7, 20, 21, 22, 23] or those that require the use of external textual information [24].
+Note that we omit subgraph methods that have been to shown to prohibitively expensive [7, 20, 21, 22, 23] or those that require the use of external textual information [24, 25].
 Personalized PageRank: PageRank [12] computes the probability of finishing a random walk of arbitrary length at some node $u$, when there is equal probability of beginning the walk at any node in the graph.
@@
 where $W$ is the random walk matrix and $x_{s}$ is a one-hot vector at node $s$.
-Intuitively, we can see that weight given to a walk decays with the increase in length due to $(1-\alpha)^{k}$.
+Observe that the weight given to a walk decays with the increase in length due to $(1-\alpha)^{k}$.
 As such, the PPR score will often be higher for those nodes of shorter distance to $s$.
-Note that for a KG, we obtain the PPR matrix by first converting the inference graph, $\mathcal{G}_{\text{inf}}$, to an undirected graph.
-We note that this is common practice in KGC [26] whereby inverse relations are added to the graph.
+For a KG, we obtain the PPR matrix by first converting the inference graph, $\mathcal{G}_{\text{inf}}$, to an undirected graph.
+This is common practice in KGC [27] whereby inverse relations are added to the graph.
 We further assume an edge weight of 1 for all edges.
+As such, the relations are completely disregarded when computing the PPR.
 3 Preliminary Study
@@
 3.1 Personalized Pagerank (PPR) is a Strong Baseline for Inductive KGC
+(a) % Increase in PPR performance from parent to inductive dataset
+(b) $\Delta\text{SPD}$ vs. PPR Hits@10
+Figure 2: (a) Percent increase in PPR performance on the inductive datasets as compared to their parent transductive dataset.
+For datasets with multiple splits (i.e., FB and WN) we take the mean performance across each.
+(b) Relationship between $\Delta\text{SPD}$ and performance.
+We observe a strong relationship among both transductive and inductive datasets.
 We begin by obtaining the PPR matrix of the inference graph for use in evaluation.
@@
 Given this new graph, for a query $(s,r,*)$, we calculate the PPR score for all possible entities $o\in\mathcal{V}$.
-Using these scores, we can obtain the rank of the true entity for the query and compute the performance in terms of Hits@10.
+Using these scores, we can obtain the rank of the true entity for our query.
+We emphasize that PPR is (a) a non-learnable heuristic, (b) ignores the relations in the KG, and (c) has no basis in KG literature as a method to perform KGC.
 In Figure 1, we show the performance when using the PPR versus the SOTA performance among supervised methods.
+The SOTA method is dataset dependent, with the specific methods listed in Appendix A.1.
 We split the datasets by transductive, (E), and (E, R) inductive.
 We include those datasets most often used in each task, comprising 25 in total.
-See Appendix B.1 for more details on the datasets chosen.
-The full set of results can be found in Appendix A.1.
+Please see Appendix B.1 for more on the datasets chosen.
+The full set of results can also be found in Appendix A.1.
 We observe that on both types of inductive datasets, the PPR score does reasonably well, performing on average only 25-29% less than SOTA.
-Also, on some datasets like the WN or ILPC inductive splits, the performance nearly matches or exceeds the supervised performance.
+This is surprising as PPR is both non-learnable and ignores the relational aspect of the KG.
+We also find that on some datasets like the WN or ILPC inductive splits, the performance nearly matches or exceeds the supervised performance.
 On the other hand, for the transductive datasets, the performance disparity is often much larger.
-Interestingly, PPR still performs well on some transductive datasets, including the popular WN18RR.
-This tells us that this phenomenon is not unique to inductive datasets.
-However, this phenomenon is far less common on transductive datasets.
-Curiously, for the inductive datasets, their performance using PPR is much higher than their transductive parents.
-For example, on the inductive datasets derived from FB15k-237, the performance of PPR increases by an average of 1,186%.
-Furthermore, for those derived from WN18RR, they see an average increase of 43%.
+Interestingly, we note that PPR still performs well on some transductive datasets, including the popular WN18RR.
+This tells us that this phenomenon is not necessarily unique to inductive datasets, but is most apparent there.
+Furthermore, we find that for the inductive datasets, their PPR performance is much higher than their transductive parents.
+We detail this in Figure 2(a) where for different inductive datasets we see the percent increase in PPR performance from transductive to inductive.
+For example, on FB15k-237 the PPR Hits@10 is 2.7%, however the mean performance on the four FB (E) splits is 42.7%, representing a 1481% increase.
+We can see that the % increase on each inductive dataset is large, with the smallest being 42% on WN.
 This suggests that there is some change in the underlying inductive graphs that are causing the performance of PPR to increase.
+Lastly, we further explore other potential shortcuts in Appendix A.3, finding that PPR is by far the most severe.
 3.2 When Does PPR Perform Well and Why?
@@
 This raises the question – why can PPR perform well on some datasets but not others?
-Figure 2: $\Delta\text{SPD}$ vs. PPR Hits@10.
 We find that the answer lies in comparing the mean shortest path distance (SPD) between positive and negative samples.
-For a query $(s,r,*)$, let $o^{+}$ denote the correct answer and $o^{-}\in\mathcal{V}^{-}(s,r)$ as the set of negative answers for the query.
-We compute the SPD on the inference graph for both $o^{+}$ and $\mathcal{V}^{-}(s,r)$.
+For a query $(s,r,*)$, let’s denote $o^{+}$ as the correct answer and $o^{-}\in\mathcal{V}^{-}(s,r)$ as the set of negative answers for the query.
+We compute the SPD on the inference graph for both the $o^{+}$ and $\mathcal{V}^{-}(s,r)$.
 This is repeated for each test sample $(s,q,*)\in\mathcal{E}_{\text{inf}}$.
@@
 This tells us, on average, how much shorter the distance between entities in positive samples are relative to those in negatives.
-In Figure 2 we show the relationship between $\Delta\text{SPD}$ and the Hits@10 when using PPR.
+In Figure 2(b) we show the relationship between $\Delta\text{SPD}$ and the Hits@10 when using PPR.
 The results are across 25 different datasets, for both inductive and transductive KGC.
@@
 The larger the discrepancy between the mean distances, the better PPR can perform.
-This suggests that the PPR scores can exploit this distance disparity in the datasets to achieve good performance, even while completely ignoring the relational information.
+This suggests that the PPR scores can exploit this pattern in the datasets, to achieve good performance, even while completely ignoring the relational information.
 But, why can PPR exploit this pattern?
 As shown in Eq.
-(1), the PPR score between two nodes is the weighted sum of random walk probabilities between them.
+(1), the PPR score between two nodes is the weighted sum of walks between them.
 Furthermore, walks of shorter length are weighed more heavily than those of longer length.
 Therefore, nodes with a shorter distance between them will be able to benefit from these highly-weighted walks, while those of larger distance will not.
-For example, when $\alpha=0.15$, the weight of a walk with $\text{SPD}=2$ is $0.72$ while when $\text{SPD}=5$ it is $0.44$.
+For example, when $\alpha=0.15$, the highest weight for a walk when $\text{SPD}=2$ is $0.72$ while when $\text{SPD}=5$ it is $0.44$.
 As such, the PPR score will invariably favor those node pairs with a lower SPD.
@@
 However, one remaining question is – why is this trend so pervasive on inductive datasets but rarely on their transductive counterparts?
-For example, on FB15k-237 [28], a transductive dataset, the Hits@10 of PPR is 2.7%.
+For example, on FB15k-237, a transductive dataset, the Hits@10 of PPR is 2.7%.
 However, across eight different inductive versions of FB15k-237, the mean PPR performance is 32%.
@@
 Since the gap between the distances is almost always larger than the original transductive dataset, this shortcut becomes more pronounced, thereby leading to a better PPR performance (see Appendix A.1 for detailed results).
-But why does the mean SPD drastically change for negative samples but not positive?
-We find that this is due to the fact that entities in positive samples are more well-connected those those in negatives.
-In Table 1, we show the % difference in mean PPR score between positive and negative samples when controlling for the SPD on two inductive datasets.
-Table 1: % Increase in in Pos over Neg PPR, controlling for SPD.
-SPD Range WN18RR v4 FB15k-237 v4
+Table 1: % Increase in Positive vs. Negative sample PPR, broken down by SPD.
+SPD WN18RR v4 FB15k-237 v4
 $[1,2)$ +17% +5%
@@
 $[4,\infty)$ +2837% +44%
+But why does the mean SPD drastically change for negative samples but not positive?
+We find that entities in positive samples are more well-connected those those in negatives.
+In Table 1, we show the % difference in PPR score between positive and negative samples when controlling for the SPD on two inductive datasets.
 We can see that the PPR score is much greater for positive samples, even for higher values of SPD.
 Therefore, the SPD of positive samples are better able to “withstand" changes in the underlying graph better than negatives, as they typically contain additional shorter walks between samples.
-However, negatives samples are typically much less well-connected, so they are more affected by removing a portion of edges from the original graph.
-(a) $\Delta\text{SPD}$ vs. PPR Hit@10
-(b) PPR Hit@10 vs. Train Size
-(c) PPR Hit@10 vs. Inference Size
-Figure 4: Experiments when generating inductive datasets from FB15k-237.
-(a) The generated datasets always have a much higher $\Delta\text{SPD}$ than the transductive dataset.
-(b)/(c) There is a trade-off between the size of the train and inference graph, making it difficult to obtain a good split.
+However, negatives are typically much less well-connected, so they are more affected by removing a portion of edges from the original graph.
 Observation 2: Constructing a good inductive dataset is difficult.
@@
 This suggests that a new strategy is needed for sampling graphs for inductive KGC from transductive datasets.
+(a) $\Delta\text{SPD}$ vs. PPR Hit@10
+(b) Hits@10 vs. Train Size
+(c) Hits@10 vs. Inference Size
+Figure 4: Experiments when generating inductive datasets from FB15k-237.
+(a) The generated datasets always have a much higher $\Delta\text{SPD}$ than the transductive dataset.
+(b)/(c) There is a trade-off between the size of the train and inference graph, making it difficult to obtain a good split.
 4 Constructing Inductive Datasets through Graph Partitioning
@@
 We note that the task of constructing inductive datasets from an existing graph $\mathcal{G}$ can be framed as a graph partitioning problem.
-Formally, we want to sample two non-overlapping partitions from the graph such that $\mathcal{G}_{\text{train}},\mathcal{G}_{\text{inf}}\in\mathcal{G}$.
+Formally, we want to sample two non-overlapping partitions from the graph such that $\mathcal{G}_{\text{train}},\mathcal{G}_{\text{inf}}\subset\mathcal{G}$ and $\mathcal{G}_{\text{train}}\cap\mathcal{G}_{\text{inf}}=\emptyset$.
 Given the analysis in Section 3.2, we hope to sample subgraphs such that $\Delta\text{SPD}(\mathcal{G}_{\text{train}})\approx\Delta\text{SPD}(\mathcal{G}_{\text{inf}})\approx\Delta\text{SPD}(\mathcal{G}$).
@@
 This is because there already exists little relationship between the two partitions in the original graph.
-Table 2: $\Delta\text{SPD}$ and PPR Performance of New and Old Inductive Datasets vs. Transductive Parents.
-Task Dataset PPR Hits@10 $\Delta\text{SPD}$
-Trans. Ind. New Ind. Old Trans. Ind. New Ind. Old
-(E) WN18RR 46.2 45.1 66.0 4.1 3.2 6.3
-CoDEx-M 9.0 11.2 21.1 0.2 0.24 0.78
-HetioNet 2.7 2.4 NA 0.55 0.29 NA
-(E, R) FB15k-237 2.7 10.8 21.4 0.46 0.48 2.42
-CoDEx-M 9.0 13.2 NA 0.20 0.42 NA
 Figure 5: Sampling Two Subgraphs
@@
 See Section B.2 for more details on the dataset creation process.
+Table 2: Mean $\Delta\text{SPD}$ and PPR Hits@10 of new and old Inductive (“Ind.”) datasets vs.
+their Transductive (“Trans.”) parent.
+We highlight the value of the inductive dataset closer to the transductive as blue.
+For every case, the $\Delta\text{SPD}$ and PPR Hits@10 of the new inductive datasets are more aligned with the transductive dataset than the older inductive datasets.
+Task Dataset Mean PPR Hits@10 Mean $\Delta\text{SPD}$
+Trans. Ind. Old Ind. New Trans. Ind. Old Ind. New
+(E) WN18RR 46.2 66.0 45.1 4.1 6.3 3.2
+CoDEx-M 9.0 21.1 11.2 0.2 0.78 0.24
+HetioNet 2.4 NA 2.7 0.55 NA 0.29
+(E, R) FB15k-237 2.7 21.4 10.8 0.46 2.42 0.48
+CoDEx-M 9.0 NA 13.2 0.20 NA 0.42
 4.2 Analysis of New Datasets
@@
 For the (E, R) datasets, we use CoDEx-M [31] and FB15k-237 [33].
-Each dataset has two separate inference graph except for CoDEx-M on the (E) task which has one.
 Note that some datasets are only suitable for one task or another.
 For example, WN18RR and HetioNet have very few relations, making it nearly impossible to sample two partitions with significantly different relations for the (E, R) task.
-On the other hand, FB15k-237 contains too many relations to sample multiple graphs for the (E) task without removing many edges from either the train or inference graph.
+On the other hand, FB15k-237 contains too many relations to sample multiple graphs for the (E) tasks without removing many edges from either the train or inference graph.
 In Table 2 we show the PPR Hits@10 and $\Delta\text{SPD}$ for the inductive datasets and their original transductive dataset.
-We note that when multiple inference graphs exist, we take the mean across each inference graph.
+When multiple inference graphs exist, we take the mean across each inference graph.
 When possible, we also include a comparison against those inductive datasets that already exist.
@@
 As compared to the old inductive datasets, the PPR performance for the new inductive datasets is much lower.
-On average, the performance of PPR is 78% lower for the new inductive datasets.
+We find that the average PPR performance is 78% lower on the new inductive datasets as compared to the older datasets.
 Furthermore, the PPR performance of the new inductive datasets is very similar to the performance on the original transductive dataset.
 A similar trend can be found when comparing the $\Delta\text{SPD}$.
-This analysis shows that newer sampling procedure can indeed sample inductive datasets that are much more similar to the original transductive graph.
-Furthermore, it greatly helps mitigate the PPR shortcut.
+This analysis shows that newer sampling procedure can indeed sample inductive datasets that are much more similar to the original transductive graph, thereby greatly mitigating the PPR shortcut.
 5 Experiments
@@
 The statistics for each dataset can be found in Appendix B.2.
-Baseline Methods: We consider prominent KGC methods including NBFNet [8], RED-GNN [11], NodePiece [19], and InGram [9]. We also consider the recent foundation model ULTRA [33].
-Hyperparameters: The exact hyperparameters and their ranges differ by method due to difference in implementations.
-For each method, we tried to follow the recommended ranges used by the authors.
-For InGram [9] we tuned the learning rate in $\{1e^{-3},5e^{-4}\}$ and the number of entity layers in $\{2,4\}$.
-For NodePiece [19] we tuned the learning rate from $\{1e^{-3},1e^{-4}\}$ and the loss margin from $\{15,25\}$.
-For RED-GNN [11], we tuned the learning rate from $\{5e^{-3},5e^{-4}\}$ and the dropout from $\{0.1,0.3\}$.
-In their experiments, NBFNet [8] uses the same hyperparameters across all datasets.
-We therefore do the same.
+Baseline Methods: We consider prominent KGC methods including NBFNet [8], RED-GNN [11], NodePiece [19], InGram [9], and Neural LP [13].
+We also consider the recent foundation model ULTRA [34].
+We omit methods that have been shown to either be prohibitively expensive (e.g.,  [7, 20, 21]) or require the use of textual information [24, 25].
 The full set of experimental settings can be found in Appendix D.
 5.2 Results
+Main Results: The main results for the supervised methods can be found in Tables 3 and 4.
+We observe that on nearly every dataset, NBFNet and RED-GNN are the two best models.
+This indicates that conditional MPNNs [15], the class of model in which both belong to, are necessary for strong performance in inductive KGC.
+Interestingly, we observe that InGram struggles in the (E, R) setting.
+This is even true when the % of new relations is high.
+This runs counter to the results on older inductive datasets [9] where InGram excelled over NBFNet and RED-GNN.
+Lastly, we observe that Neural LP sometimes fails to converge, resulting in a near zero performance and thus causing the model to have a high performance variance.
 Table 3: (E) Inductive Results (Hits@10) for supervised methods.
@@
 PPR 11.2 66.2 24 3.2 2.2
+Neural LP 13.0 ± 17.9 37.9 ± 1.4 14.8 ± 1.9 12.0 ± 16.4 10.7 ± 15.0
 NodePiece 6.8 ± 0.8 29.6 ± 0.8 4.8 ± 0.6 10.2 ± 0.9 15.4 ± 0.9
@@
 Table 4: (E, R) Inductive Results for supervised methods. The % of new relations are in parentheses.
-Models FB15k-237 CoDEx-M Inference 1 (27%) Inference 2 (63%) Inference 1 (10%) Inference 2 (57%) MRR Hits@10 MRR Hits@10 MRR Hits@10 MRR Hits@10 PPR 3.8 9.1 4.7 12.4 3.3 10.9 7.1 15.4 NodePiece 1.4 ± 0.1 3.0 ± 0.6 2.6 ± 0.2 4.7 ± 0.5 1.5 ± 0.2 3.1 ± 0.6 1.6 ± 0.4 2.5 ± 1.0 InGram 11.2 ± 1.9 23.8 ± 3.0 9.2 ± 0.9 20.2 ± 2.0 10.3 ± 1.2 20.4 ± 3.3 7.2 ± 4.2 15.9 ± 10.0 RED-GNN 11.8 ± 2.8 21.6 ± 5.8 19.8 ± 2.2 33.3 ± 4.2 16.8 ± 1.6 29.2 ± 2.9 14.7 ± 5.3 26.5 ± 10.4 NBFNet 15.8 ± 0.9 27.5 ± 1.8 15.9 ± 0.2 26.2 ± 0.3 27.9 ± 8.8 47.7 ± 11.8 8.5 ± 4.8 17.6 ± 10.0
-Main Results: The main results for supervised methods can be found in Tables 3 and 4.
-We observe that on nearly every dataset, NBFNet and RED-GNN are the two best models.
-This indicates that conditional MPNNs [15], the class of model in which both belong to, are necessary for strong performance in inductive KGC.
-Interestingly, we observe that InGram struggles in the (E, R) setting.
-This is even true when the % of new relations is high.
-This run counters to the results on older inductive datasets [9] where InGram excelled over NBFNet and RED-GNN.
-Lastly, outside of NodePiece, each method is usually able to significantly outperform PPR.
-Performance Comparison on New and Old Inductive Datasets: We compare the change in performance from the previous inductive datasets to our new datasets.
-Note that we only compare when there is a corresponding dataset sampled from the same transductive dataset and for the same task.
-For the (E) task, this includes CoDEx-M (ILPC-S [10]) and WN18RR (v1-v4 [7]).
-For the (E, R) task it is only FB15k-237 (25-100 [9]).
-For each, we compare the performance of the SOTA on the older dataset to their corresponding performance on the new datasets.
-This is NBFNet [8] for WN18RR and FB15k-237 and NodePiece [19] for CoDEx-M.
-We only compare the performance of the old SOTA as not all methods were tested on the older datasets or guranteed to be well-tuned.
+Models FB15k-237 CoDEx-M
+Inference 1 (27%) Inference 2 (63%) Inference 1 (10%) Inference 2 (57%)
+PPR 9.1 12.4 10.9 15.4
+Neural LP 17.5 ± 9.9 22.4 ± 12.8 16.7 ± 22.8 9.8 ± 13.7
+NodePiece 3.0 ± 0.6 4.7 ± 0.5 3.1 ± 0.6 2.5 ± 1.0
+InGram 23.8 ± 3.0 20.2 ± 2.0 20.4 ± 3.3 15.9 ± 10.0
+RED-GNN 21.6 ± 5.8 33.3 ± 4.2 29.2 ± 2.9 26.5 ± 10.4
+NBFNet 27.5 ± 1.8 26.2 ± 0.3 47.7 ± 11.8 17.6 ± 10.0
+Performance Comparison of PPR vs.
+SOTA: In our original analysis, we showed that for the older inductive datasets, the performance gap between the SOTA method of each dataset and PPR is quite small.
+Specifically, the mean % difference between the two was relatively small at 25-29% (see Figure 1).
+We now perform the same analysis on the new inductive datasets.
 The results are shown in Figure 6(a).
-We find that performance on each of the new datasets drops significantly.
-This suggests that removing the shortcut has a large negative effect on the performance.
+We observe that PPR generally performs much worse than the SOTA method with a mean % difference of 99%.
+Furthermore, this is more in line with the transductive datasets which have a mean % difference of 121%.
+This suggests that in addition to the decrease in raw PPR performance, the relative performance of PPR also decreases on the newer datasets.
+Table 5: Mean performance on Old vs. New inductive datasets by Method. Red indicates a decrease in the mean performance on the newer splits.
+Models WN18RR (E) FB15k-237 (E, R)
+Old Ind. New Ind. % Difference Old Ind. New Ind. % Difference
+PPR 66.0 46.2 -30% 21.4 10.8 -49.5%
+Neural LP 67.6 26.4 -61.0% 16.2 20.0 +23.1%
+NodePiece 29.8 17.2 -42.3% 5.0 3.9 -23.0%
+InGram 49.9 23.0 -53.9% 29.6 22.0 -25.7%
+RED-GNN 70.6 50.3 -28.8% 25.0 27.5 +9.8%
+NBFNet 72.2 52.5 -27.4% 24.8 26.9 +8.3%
+Mean - - -40.6% - - -9.5%
+Performance Comparison on New and Old Inductive Datasets: Previously, in Table 2, we compared both the PPR performance and $\Delta\text{SPD}$ on the new and older inductive datasets.
+We found that both metrics tend to be much higher on the older inductive datasets, indicating that our newer splits are effective in mitigating the shortcut.
+Given those results, a natural question is whether we see a similar drop in performance for neural methods?
+We limit our analysis to WN18RR (E) and FB15k-237 (E, R).
+This is either due to a lack of older datasets (i.e., CoDEX-M (E, R) and HetioNet (E)) or minimal results on the older datasets, i.e., CoDEX-M (E).
+For each dataset, we compute the mean performance across inference graphs.
+The results are shown in Table 5.
+We find that performance of most methods drops significantly across both datasets, with an average drop of 40.6% and 9.5% on WN18RR (E) and FB15k-237 (E, R), respectively.
+This suggests that removing the shortcut has a large negative effect on the performance, thereby suggesting that our new datasets are indeed harder.
 Performance of ULTRA [34]: We further compare against ULTRA, a recent foundation model designed for fully inductive KGC.
 We evaluate ULTRA under the 0-shot setting.
-See Appendix E for more details on the version of ULTRA used.
+See Appendix E for more details on the versions of ULTRA used.
 The results are in Figure 6(b) where for each dataset, we average the results across the different inference graphs (full results in Appendix A.2).
@@
 However, on the (E, R) task, ULTRA significantly outperforms other methods.
-This suggests that ULTRA contains a much greater ability to generalize to unseen relations than existing supervised methods.
-(a) Performance of SOTA on Old vs. New Datasets
-(b) Results across Inf. graphs for ULTRA vs. SOTA
-Figure 6: (a) Performance of SOTA on old vs. new datasets, (b) ULTRA [33] vs. supervised SOTA.
+This suggests that ULTRA contains a much greater ability than other existing methods to generalize to unseen relations.
+(a) PPR vs. SOTA Hits@10 on newer datasets
+(b) Mean Hits@10 for ULTRA vs. SOTA
+Figure 6: (a) Performance of PPR vs.
+SOTA on new inductive datasets.
+Note that we abbreviate CoDEx-M as CM, FB15k-237 as FB, HetioNet as HN and WN18RR as WN.
+Furthermore, the number represents the inference graph.
+(b) Mean performance of the foundation model ULTRA [34] vs.
+the supervised SOTA on the newer datasets.
 6 Conclusion
@@
 Through our study, we uncover that this shortcut is due to how inductive datasets are created.
-To remedy this problem, we propose a new dataset construction process based on graph partitioning that helps mitigate the impact of the studied shortcut.
+To remedy this problem, we propose a new dataset construction process based on graph partitioning that empirically mitigates the impact of the studied shortcut.
 We then construct new benchmark datasets using this new procedure and benchmark various methods.
+Examining the results, we observe that the relative performance decreases on most datasets and methods, indicating that the newer benchmark datasets are harder than the previous ones.
+For future work, we plan to explore creating inductive KG datasets that aren’t sampled from existing transductive datasets.
 Appendix A Additional Results
@@
 We further include the % difference in performance and which method is considered SOTA.
-These are shown in Tables 56, and 7 for the (E), (E, R), and transductive datasets, respectively.
+These are shown in Tables 67, and 8 for the (E), (E, R), and transductive datasets, respectively.
 For an overview of SOTA performance on additional KG datasets, please see [34].
@@
 Hits@10 69.6 64.5 45.7 91.3
+A.3 Performance of Other Potential Shortcuts
+In our study, we show that PPR can achieve strong performance on most inductive KG datasets, indicating a shortcut.
+A natural question is whether this is true for just PPR, or if other shortcuts exist.
+Another bias discussed in KG literature is degree bias [38].
+In their study, they show that that KG methods tend to perform better on entities with a higher degree.
+Specifically, they observe that what matters is the degree of the entity being predicted (they refer to this as the “tail degree”).
+Based on this, we study whether the tail degree can is a good predictor of performance.
+We show that results on a number of inductive and transductive datasets in Table 11.
+We report Hits@10 for all results.
+We find that the tail degree often achieves a much lower performance than PPR.
+Specifically, for the datasets in Table  11, while the PPR performance is on average only 31% lower than the SOTA, the tail degree performance is 84% lower.
+This indicates the severity of the PPR shortcut is much more severe as compared to other known biases like the tail degree.
+Table 11: Performance of Tail Degree [38], PPR, and SOTA method. Performance is in terms of Hits@10. We further bold the larger of the PPR and tail degree performance.
+Task Dataset Tail Degree PPR SOTA
+Inductive WN v1 10.4 77.1 82.6
+WN v2 5.1 74.4 83.6
+WN v3 10.1 45.2 58.2
+WN v4 2.2 67.3 74.5
+FB v1 20.5 41.2 60.7
+FB v2 22.7 47.6 70.4
+FB v3 16.8 43.5 67.7
+FB v4 16.1 38.4 66.8
+Transductive WN18RR 2.0 46.2 59.9
+FB15k-237 6.0 2.7 66.6
+DBPedia100k 2.9 30.2 41.8
 Appendix B Datasets
 B.1 Existing Datasets
-We detail the statistics of all existing transductive and inductive datasets in Tables 10, 11, and 12, respectively.
+We detail the statistics of all existing transductive and inductive datasets in Tables 12, 13, and 14, respectively.
 We further include the licenses for each in Table 15.
@@
 B.2 New Datasets
-The raw data for the new datasets introduced in Section 4, can be found at https://github.com/HarryShomer/Better-Inductive-KGC/tree/master/new_data.
-Please see the README for more information on how to load the data.
-We further detail the statistics of all the new datasets in Tables 14 and 15.
+We further detail the statistics of all the new datasets in Tables 16 and 17.
 Table 16: Statistics for New (E) Inductive Datasets.
@@
 All models were implemented in PyTorch [41] and Torch-Geometric [42].
+Hyperparameters: Due to the different types and ranges of hyperparameters used for each method, the exact hyperparameters and their ranges differ by method.
+For each method, we tried to follow the recommended ranges used by the authors.
+For InGram [9] we tuned the learning rate in $\{1e^{-3},5e^{-4}\}$ and the number of entity layers in $\{2,4\}$.
+For NodePiece [19] we tuned the learning rate from $\{1e^{-3},1e^{-4}\}$ and the loss margin from $\{15,25\}$.
+For RED-GNN [11], we tuned the learning rate from $\{5e^{-3},5e^{-4}\}$ and the dropout from $\{0.1,0.3\}$.
+In their experiments, NBFNet [8] uses the same hyperparameters across all datasets.
+We therefore do the same.
+For Neural LP [13] we also use the default set of hyperparameters that is shared across datasets.
 Evaluation: For a given test triple $(s,r,o)$, we strive to predict both $s$ and $o$ individually.
@@
 As such, following [4] we use the filtered setting and remove those entities from the ranking.
-Once we obtain the rank of the true entity (i.e., “how many negative entities does it have a higher probability than”) we compute two metrics – the mean reciprocal rank (MRR) and Hits@K.
-MRR is given by $1/\text{Rank}$ such that when Eq.
-2 holds true for $\forall v\in\mathcal{V}$ it will be equal to 1.
-Hits@K computes whether $\text{Rank}\leq K$.
+Once we obtain the rank of the true entity (i.e., “how many negative entities does it have a higher probability than”) we compute the Hits@K%.
+Hits@Kcomputes the percentage of sample where the following inequality holds true, $\text{Rank}\leq K$.
 Appendix E ULTRA Settings
@@
 Dataset Inference Graph W/o With % Difference
-CoDEx-M (E) 0 46.6 50.2 -7.2%
-WN18RR (E) 0 71.9 72.7 +1.1%
+CoDEx-M (E) 1 46.6 50.2 -7.2%
+WN18RR (E) 1 71.9 72.7 +1.1%
 2 38.8 38.5 -0.8%
-FB15k-237 (E, R) 0 69.6 72.1 -4.5%
+FB15k-237 (E, R) 1 69.6 72.1 -4.5%
 2 64.5 64.8 -0.5%
-CoDEx-M (E) 0 45.7 50.4 -9.3%
+CoDEx-M (E, R) 1 45.7 50.4 -9.3%
 2 91.1 91.3 -0.2%
 Appendix F Implementation
-All code used for experiments in this study can be found in the following repository – https://github.com/HarryShomer/Better-Inductive-KGC.
+All code used for experiments in this study can be found in the following anonymized repository – https://github.com/HarryShomer/Better-Inductive-KGC.
 Please see the README in the repository for more information on how to run the different experiments.

### v3

@@
 Towards Better Benchmark Datasets for Inductive Knowledge Graph Completion
-Abstract
+Abstract.
 Knowledge Graph Completion (KGC) attempts to predict missing facts in a Knowledge Graph (KG).
@@
 However, we find that the current procedure for constructing inductive KGC datasets inadvertently creates a shortcut that can be exploited even while disregarding the relational information.
-Specifically, we observe that the Personalized PageRank (PPR) score can achieve strong or near SOTA performance on most inductive datasets.
+Specifically, we observe that the Personalized PageRank (PPR) score can achieve strong or near SOTA performance on most datasets.
 In this paper, we study the root cause of this problem.
@@
 The new benchmark datasets help promote a better understanding of the capabilities and challenges of inductive KGC by removing any shortcuts that obfuscate performance.
-1 Introduction
+The code and dataset and can be found https://github.com/HarryShomer/Better-Inductive-KGC.
+Knowledge graphs, Link prediction
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06journalyear: 2025copyright: ccconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3737438isbn: 979-8-4007-1454-2/2025/08ccs: Computing methodologies Machine learning
+1. Introduction
 Knowledge Graph Completion (KGC) attempts to predict unseen facts given an existing knowledge graph (KG).
-KGC has numerous applications including drug discovery [1], personalized medicine [2], and recommendations [3].
+KGC has numerous applications including drug discovery (Zeng et al., 2022), personalized medicine (Chandak et al., 2023), and recommendations (Wang et al., 2021).
 Traditionally, most research on KGC was focused on the transductive setting, where the same sets of entities and relations are seen during training and testing.
-Most methods [4, 5, 6] generally focus on learning embeddings for all entities and relations to facilitate the prediction of new facts.
+Most methods (Bordes et al., 2013; Trouillon et al., 2016; Schlichtkrull et al., 2018) generally focus on learning embeddings for all entities and relations to facilitate the prediction of new facts.
 In recent years, interest in KGC has shifted towards designing methods that can generalize to new entities or relations not seen during training.
 This task, known as “inductive KGC”, requires a method to train on a graph $\mathcal{G}_{\text{train}}$ and perform inference on a different graph $\mathcal{G}_{\text{inf}}$, where the inference graph contains either new entities, relations, or both.
-Because of this, methods for inductive KGC don’t rely on fixed embeddings for entities or relations, instead opting for more flexible techniques that can inductively learn representations based on a given graph [7, 8, 9].
-To asses the ability of methods for this task, new datasets have been constructed that require methods to reason inductively.
-All inductive datasets [7, 10, 9] are constructed from existing transductive KGC datasets.
+Because of this, methods for inductive KGC don’t rely on fixed embeddings for entities or relations, instead opting for more flexible techniques that can inductively learn representations based on a given graph (Teru et al., 2020; Zhu et al., 2021; Lee et al., 2023).
+To assess the ability of methods for this task, new datasets have been constructed that require methods to reason inductively.
+All inductive datasets (Teru et al., 2020; Galkin et al., 2022; Lee et al., 2023) are constructed from existing transductive KGC datasets.
 This is done by sampling two graphs, one each for train and inference, which contain disjoint entities.
-Multiple methods [8, 11, 9] have reported tremendous promise on these newer benchmark datasets.
-However, we find that on almost all inductive datasets, we can achieve competitive performance by using the Personalized PageRank [12] (PPR) score to perform inference.
+Multiple methods (Zhu et al., 2021; Zhang and Yao, 2022; Lee et al., 2023) have reported tremendous promise on these newer benchmark datasets.
+However, we find that on almost all inductive datasets, we can achieve competitive performance by using the Personalized PageRank (PPR) score (Page et al., 1999) to perform inference.
 We note that PPR is a non-learnable heuristic and ignores the relational information in the graph.
@@
 (c) Transductive
-Figure 1: Hits@10 of PPR vs. Supervised SOTA. Results are on (a) (E) Inductive, (b) (E, R) Inductive, and (c) Transductive datasets.
-2 Background and Related Work
+Figure 1. Hits@10 of PPR vs. Supervised SOTA. Results are on (a) (E) Inductive, (b) (E, R) Inductive, and (c) Transductive datasets.
+2. Background and Related Work
 Throughout this study we denote a knowledge graph as $\mathcal{G}=\{\mathcal{V},\mathcal{R},\mathcal{E}\}$ where $\mathcal{V}$ are the set of entities (i.e., nodes), $\mathcal{R}$ the set of relations (i.e., edge types), and $\mathcal{E}$ the set of edges (i.e., triples) of the form $(s,r,o)$ where $s$ and $o$ are entities and $r$ a relation.
@@
 All existing inductive datasets are sampled from existing transductive datasets.
-Furthermore, the majority of inductive datasets [7, 9] are further created via the same procedure introduced by [7].
+Furthermore, the majority of inductive datasets (Teru et al., 2020; Lee et al., 2023) are further created via the same procedure introduced by (Teru et al., 2020).
 We now give a brief overview of this procedure.
@@
 The inference graph is then sampled in a similar manner from the resulting graph $\mathcal{G}\setminus\mathcal{G}_{\text{train}}$.
-One exception to this procedure are the ILPC datasets, introduced by [10], which instead sample $p\%$ of the nodes from $\mathcal{G}$ and use them to create $\mathcal{G}_{\text{train}}$.
+One exception to this procedure are the ILPC datasets, introduced by (Galkin et al., 2022), which instead sample $p\%$ of the nodes from $\mathcal{G}$ and use them to create $\mathcal{G}_{\text{train}}$.
 The rest of the nodes are then used to construct $\mathcal{G}_{\text{inf}}$.
 In practice, we find that both methods tend to produce similar graphs.
-Inductive KGC Methods: NeurlLP [13] and DRUM [14] consider combining the path representations of different length between both entities in a triple.
+Inductive KGC Methods: NeurlLP (Yang et al., 2017) and DRUM (Sadeghian et al., 2019) consider combining the path representations of different length between both entities in a triple.
 However, since they explicitly consider each path, they are often limited to only considering paths of up to length 2 or 3.
-Conditional MPNNs [15] are a more efficient alternative to encoding higher-order path information.
+Conditional MPNNs (Huang et al., 2024) are a more efficient alternative to encoding higher-order path information.
 They work by conditioning the message passing mechanism on the known entity, allowing the implicit aggregation of all paths up to a length $L$ (which is equal to the number of GNN layers).
-The value of $L$ is typically $5/6$.
-Prominent examples include NBFNet [8] and RED-GNN [11].
-More scalable alternatives have been proposed that prune the propagated messages [16, 17, 18].
-NodePiece [19] is concerned with parameter efficiency, using an anchor-based approach to learn a more compact set of entity and relation representations.
+The value of $L$ is typically $5$ or $6$.
+Prominent examples include NBFNet (Zhu et al., 2021) and RED-GNN (Zhang and Yao, 2022).
+More scalable alternatives have been proposed that prune the propagated messages (Zhu et al., 2023; Zhang et al., 2023; Shomer et al., 2023b).
+NodePiece (Galkin et al., 2021) is concerned with parameter efficiency, using an anchor-based approach to learn a more compact set of entity and relation representations.
 All previous methods assume that a fixed set of relations exist between the train and inference graphs.
-To account for new relations in inference, InGRAM [9] introduces the concept of a “relation graph”, which inductively encodes the representation of each relation.
-Note that we omit subgraph methods that have been to shown to prohibitively expensive [7, 20, 21, 22, 23] or those that require the use of external textual information [24, 25].
-Personalized PageRank: PageRank [12] computes the probability of finishing a random walk of arbitrary length at some node $u$, when there is equal probability of beginning the walk at any node in the graph.
-Personalized PageRank (PPR) [12] is a version of PageRank that is “personalized" to some root node $s$, where at each step there is a probability $\alpha$ of teleporting to $s$.
-The set of PPR scores for a root node $s$ is given by $\text{pr}_{s}\in\mathbb{R}^{\lvert\mathcal{V}\rvert}$ and can be formulated as the weighted sum of all random walk probabilities between two nodes [26]:
-$\text{pr}_{s}=\alpha\sum_{k=0}^{\infty}(1-\alpha)^{k}W^{k}x_{s},$ (1)
-where $W$ is the random walk matrix and $x_{s}$ is a one-hot vector at node $s$.
+To account for new relations in inference, InGRAM (Lee et al., 2023) introduces the concept of a “relation graph”, which inductively encodes the representation of each relation.
+Gao et al.
+(2023) introduces the concept of “double permutation- equivariant representations” as a way to model KGs that are equivariant to permutations of both the entities and relations.
+They theoretically show that capturing this property is essential for proper generalization across KGs.
+To this point these introduce a new method, ISDEA/ISDEA+, that can satisfy this property.
+They further introduce a variant of InGRAM (Lee et al., 2023), DEq-InGram, that endows it with the ability to compute double equivariant representations.
+Note that we omit subgraph methods that have been to shown to prohibitively expensive (Teru et al., 2020; Liu et al., 2021; Mai et al., 2021; Xu et al., 2022; Geng et al., 2023) or those that require the use of external textual information (Gesese et al., 2022; Daza et al., 2021).
+Personalized PageRank: PageRank (Page et al., 1999) computes the probability of finishing a random walk of arbitrary length at some node $u$, when there is equal probability of beginning the walk at any node in the graph.
+Personalized PageRank (PPR) (Page et al., 1999) is a version of PageRank that is “personalized" to some root node $s$, where at each step there is a probability $\alpha$ of teleporting to $s$.
+The set of PPR scores for a root node $s$ is given by $\text{pr}_{s}\in\mathbb{R}^{\lvert\mathcal{V}\rvert}$ and can be formulated as the weighted sum of all random walk probabilities between two nodes (Chung, 2007):
+(1) $\text{pr}_{s}=\alpha\sum_{k=0}^{\infty}(1-\alpha)^{k}W^{k}x_{s},$
+where $W=D^{-1}A$ and $x_{s}$ is a one-hot vector at node $s$.
 Observe that the weight given to a walk decays with the increase in length due to $(1-\alpha)^{k}$.
@@
 For a KG, we obtain the PPR matrix by first converting the inference graph, $\mathcal{G}_{\text{inf}}$, to an undirected graph.
-This is common practice in KGC [27] whereby inverse relations are added to the graph.
+This is common practice in KGC (Dettmers et al., 2018) whereby inverse relations are added to the graph.
 We further assume an edge weight of 1 for all edges.
 As such, the relations are completely disregarded when computing the PPR.
-3 Preliminary Study
+3. Preliminary Study
 In this section, we examine the performance on common inductive KGC datasets.
@@
 and (2) Why can PPR sometimes perform well for KGC?
-3.1 Personalized Pagerank (PPR) is a Strong Baseline for Inductive KGC
+3.1. Personalized Pagerank (PPR) is a Strong Baseline for Inductive KGC
 (a) % Increase in PPR performance from parent to inductive dataset
-(b) $\Delta\text{SPD}$ vs. PPR Hits@10
-Figure 2: (a) Percent increase in PPR performance on the inductive datasets as compared to their parent transductive dataset.
+(b) $\Delta\text{SPD}$ vs. PPR Hits@10 for all datasets
+Figure 2.
+(a) Percent increase in PPR performance on the inductive datasets as compared to their parent transductive dataset.
 For datasets with multiple splits (i.e., FB and WN) we take the mean performance across each.
@@
 In Figure 1, we show the performance when using the PPR versus the SOTA performance among supervised methods.
-The SOTA method is dataset dependent, with the specific methods listed in Appendix A.1.
+The SOTA method is dataset dependent, with the specific methods are shown in Appendix A.1..
 We split the datasets by transductive, (E), and (E, R) inductive.
@@
 Please see Appendix B.1 for more on the datasets chosen.
-The full set of results can also be found in Appendix A.1.
+The full set of results are shown in Table 6.
 We observe that on both types of inductive datasets, the PPR score does reasonably well, performing on average only 25-29% less than SOTA.
-This is surprising as PPR is both non-learnable and ignores the relational aspect of the KG.
+This is surprising as PPR is both non-learnable and ignores the relations of the KG.
 We also find that on some datasets like the WN or ILPC inductive splits, the performance nearly matches or exceeds the supervised performance.
@@
 Lastly, we further explore other potential shortcuts in Appendix A.4, finding that PPR is by far the most severe.
-3.2 When Does PPR Perform Well and Why?
+3.2. When Does PPR Perform Well and Why?
 In the previous section we detailed that PPR score performs well on the inductive setting.
@@
 We then calculate the mean SPD across all positive and negative test samples, which we denote as $\text{SPD}^{+}$ and $\text{SPD}^{-}$, respectively.
-The different in mean SPD is correspondingly given by $\Delta\text{SPD}=\text{SPD}^{+}-\text{SPD}^{-}$.
+The difference in mean SPD is correspondingly given by $\Delta\text{SPD}=\text{SPD}^{-}-\text{SPD}^{+}$.
 This tells us, on average, how much shorter the distance between entities in positive samples are relative to those in negatives.
@@
 The larger the discrepancy between the mean distances, the better PPR can perform.
-This suggests that the PPR scores can exploit this pattern in the datasets, to achieve good performance, even while completely ignoring the relational information.
+This suggests that the PPR scores can exploit this pattern in the datasets, to achieve good performance, even while completely ignoring the relation information.
 But, why can PPR exploit this pattern?
@@
 As such, the PPR score will invariably favor those node pairs with a lower SPD.
-Note that we use PPR as opposed to SPD in our experiments since all-pairs SPD is costly to calculate and the full PPR matrix can be efficiently approximated via [28].
+Table 1. % Increase in Positive vs. Negative sample PPR, broken down by SPD.
+SPD WN18RR v4 FB15k-237 v4
+$[1,2)$ +17% +5%
+$[2,3)$ +29% +200%
+$[3,4)$ +82% +328%
+$[4,\infty)$ +2837% +44%
+Note that we use PPR as opposed to SPD in our experiments since all-pairs SPD is costly to calculate and the full PPR matrix can be efficiently approximated via (Andersen et al., 2006).
 Furthermore, as we show in Section 3.3, even when controlling for the SPD, the PPR score can help differentiate between positive and negative samples.
-3.3 Why is This So Common on Inductive Datasets?
+3.3. Why is This So Common on Inductive Datasets?
 (a) FB15k-237
 (b) WN18RR
-Figure 3: The mean shortest path (SP) distance between positive test samples and negative samples.
+Figure 3.
+The mean shortest path (SP) distance between positive test samples and negative samples.
 We display the results for two transductive datasets (FB15k-237 and WN18RR) and the inductive datasets that are derived from them.
@@
 This creates the shortcut described in Section 3.2, where the SPD can easily differentiate between positive and negative samples.
-Since the gap between the distances is almost always larger than the original transductive dataset, this shortcut becomes more pronounced, thereby leading to a better PPR performance (see Appendix A.1 for detailed results).
-Table 1: % Increase in Positive vs. Negative sample PPR, broken down by SPD.
-SPD WN18RR v4 FB15k-237 v4
-$[1,2)$ +17% +5%
-$[2,3)$ +29% +200%
-$[3,4)$ +82% +328%
-$[4,\infty)$ +2837% +44%
+Since the gap between the distances is almost always larger than the original transductive dataset, this shortcut becomes more pronounced, thereby leading to a better PPR performance.
 But why does the mean SPD drastically change for negative samples but not positive?
@@
 Furthermore, there is an inverse relationship between the size of the train and inference graph, making it hard to find a “sweet spot” where the PPR performance is low and both graphs aren’t too small.
+Discussion.
 The studies in this section show that the current strategy for constructing inductive datasets from existing transductive datasets is liable to introduce a well-performing shortcut into the existing graph.
@@
 (c) Hits@10 vs. Inference Size
-Figure 4: Experiments when generating inductive datasets from FB15k-237.
+Figure 4.
+Experiments when generating inductive datasets from FB15k-237.
 (a) The generated datasets always have a much higher $\Delta\text{SPD}$ than the transductive dataset.
-(b)/(c) There is a trade-off between the size of the train and inference graph, making it difficult to obtain a good split.
-4 Constructing Inductive Datasets through Graph Partitioning
+Furthermore, there is a trade-off between the size of the train (b) and inference graph (c), making it difficult to obtain a good split.
+4. Constructing Inductive Datasets through Graph Partitioning
 In the previous section, we showed that the PPR score can achieve strong performance on most inductive KGC datasets.
@@
 In the next subsection, we introduce our strategy which utilizes graph partitioning to alleviate this problem.
-4.1 Partition-Based Dataset Sampling
+4.1. Partition-Based Dataset Sampling
 We’ve previously covered in Section 3 that the existing procedure for creating inductive datasets leads to suboptimal subgraphs.
@@
 We note that the task of constructing inductive datasets from an existing graph $\mathcal{G}$ can be framed as a graph partitioning problem.
-Formally, we want to sample two non-overlapping partitions from the graph such that $\mathcal{G}_{\text{train}},\mathcal{G}_{\text{inf}}\subset\mathcal{G}$ and $\mathcal{G}_{\text{train}}\cap\mathcal{G}_{\text{inf}}=\emptyset$.
+Formally, we want to sample two non-overlapping partitions from the graph such that $\mathcal{G}_{\text{train}},\mathcal{G}_{\text{inf}}\subseteq\mathcal{G}$ and $\mathcal{G}_{\text{train}}\cap\mathcal{G}_{\text{inf}}=\emptyset$.
 Given the analysis in Section 3.2, we hope to sample subgraphs such that $\Delta\text{SPD}(\mathcal{G}_{\text{train}})\approx\Delta\text{SPD}(\mathcal{G}_{\text{inf}})\approx\Delta\text{SPD}(\mathcal{G}$).
@@
 This is because there already exists little relationship between the two partitions in the original graph.
-Figure 5: Sampling Two Subgraphs
-Multiple popular approaches [29, 30] exist that attempt to divide the graph into optimal partitions.
+Multiple popular approaches (Shi and Malik, 2000; Blondel et al., 2008) exist that attempt to divide the graph into optimal partitions.
 The guiding principle in these approaches is that the partitions should be internally dense but only sparsely connected to one another.
@@
 Also, since the partitions are created at the same time, we avoid sampling the inference graph after the training, which can be suboptimal.
-In practice, we consider using Spectral Clustering [29] or the Louvain method [30], as dependent on the dataset.
+Figure 5. Sampling Two Subgraphs
+In practice, we consider using Spectral Clustering (Shi and Malik, 2000) or the Louvain method (Blondel et al., 2008), as dependent on the dataset.
 Once the original graph is partitioned into $n$ partitions, we sample $k$ of those to be used.
@@
 Of the $k$ partitions, one is chosen as the training graph while the other $k-1$ are designated as separate inference graphs.
-This is an important advantage of using partitioning to sample the graphs, as it allows us to avoid having to sample multiple different train and inference pairs as in [7, 9], thereby allowing for more efficient benchmarking.
+This is an important advantage of using partitioning to sample the graphs, as it allows us to avoid having to sample multiple different train and inference pairs as in (Teru et al., 2020; Lee et al., 2023), thereby allowing for more efficient benchmarking.
 Lastly, we note that when sampling graphs for the (E) inductive task, we further remove new relations from the inference graph.
@@
 See Section B.2 for more details on the dataset creation process.
-Table 2: Mean $\Delta\text{SPD}$ and PPR Hits@10 of new and old Inductive (“Ind.”) datasets vs.
+Table 2.
+Mean $\Delta\text{SPD}$ and PPR Hits@10 of new and old Inductive (“Ind.”) datasets vs.
 their Transductive (“Trans.”) parent.
 We highlight the value of the inductive dataset closer to the transductive as blue.
-For every case, the $\Delta\text{SPD}$ and PPR Hits@10 of the new inductive datasets are more aligned with the transductive dataset than the older inductive datasets.
 Task Dataset Mean PPR Hits@10 Mean $\Delta\text{SPD}$
@@
 CoDEx-M 9.0 NA 13.2 0.20 NA 0.42
-4.2 Analysis of New Datasets
+4.2. Analysis of New Datasets
 In this section we analyze the new inductive datasets created following the partition-based procedure outlined in Section 4.1.
-We create (E) inductive datasets, from WN18RR [27], CoDEx-M [31], and HetioNet [32].
-For the (E, R) datasets, we use CoDEx-M [31] and FB15k-237 [33].
+We create (E) inductive datasets, from WN18RR (Dettmers et al., 2018), CoDEx-M (Safavi and Koutra, 2020), and HetioNet (Himmelstein et al., 2017).
+For the (E, R) datasets, we use CoDEx-M (Safavi and Koutra, 2020) and FB15k-237 (Toutanova and Chen, 2015).
 Note that some datasets are only suitable for one task or another.
 For example, WN18RR and HetioNet have very few relations, making it nearly impossible to sample two partitions with significantly different relations for the (E, R) task.
-On the other hand, FB15k-237 contains too many relations to sample multiple graphs for the (E) tasks without removing many edges from either the train or inference graph.
+On the other hand, FB15k-237 contains too many relations to sample multiple graphs for the (E) tasks without removing many edges from either graph.
 In Table 2 we show the PPR Hits@10 and $\Delta\text{SPD}$ for the inductive datasets and their original transductive dataset.
@@
 When possible, we also include a comparison against those inductive datasets that already exist.
-For example, for WN18RR in the (E) task, 4 datasets exist from [7].
-As compared to the old inductive datasets, the PPR performance for the new inductive datasets is much lower.
-We find that the average PPR performance is 78% lower on the new inductive datasets as compared to the older datasets.
-Furthermore, the PPR performance of the new inductive datasets is very similar to the performance on the original transductive dataset.
+For example, for WN18RR in the (E) task, 4 datasets exist from (Teru et al., 2020).
+Compared to the old inductive datasets, the PPR performance for the newer datasets is much lower.
+Specifically, the average PPR performance is 78% lower on the new inductive datasets as compared to the older datasets.
+Also, the PPR performance of the new inductive datasets is very similar to the performance on the original transductive dataset.
 A similar trend can be found when comparing the $\Delta\text{SPD}$.
-This analysis shows that newer sampling procedure can indeed sample inductive datasets that are much more similar to the original transductive graph, thereby greatly mitigating the PPR shortcut.
-5 Experiments
-5.1 Experimental Settings
+This analysis shows that newer sampling procedure can indeed sample inductive datasets that are much more similar to the original transductive graph, greatly mitigating the PPR shortcut.
+5. Experiments
+5.1. Experimental Settings
 Datasets: We use the new datasets created in Section 4.1.
@@
 For the (E, R) setting, it is FB15k-237 and CoDEx-M.
-We sample 2 inference graphs for each dataset, except for CoDEx-M on the (E) setting where we could only find 1 suitable graph for inference.
+We sample two inference graphs for each dataset, except for CoDEx-M on the (E) setting where we could only find one suitable graph for inference.
 For each inference graph, 10% of edges are randomly removed for testing.
 For validation 10% of edges are removed from the training graph.
-It is necessary that the validation samples are extracted from the train graph as the inference graphs must remain unobserved during training.
+Note that while previous studies (Teru et al., 2020; Galkin et al., 2021) validate against samples from the inference graph, it is necessary that the validation samples be extracted from the train graph.
+This is because using validation samples from the inference graph can be seen as a form of test leakage, as we are able to see the inference graph before testing.
+As such, because the inference graphs must remain unobserved during training, we can only sample validation from the training graph.
 The statistics for each dataset can be found in Appendix B.2.
-Baseline Methods: We consider prominent KGC methods including NBFNet [8], RED-GNN [11], NodePiece [19], InGram [9], and Neural LP [13].
-We also consider the recent foundation model ULTRA [34].
-We omit methods that have been shown to either be prohibitively expensive (e.g.,  [7, 20, 21]) or require the use of textual information [24, 25].
-The full set of experimental settings can be found in Appendix D.
-5.2 Results
+Evaluation: For a given test triple $(s,r,o)$, we strive to predict both $s$ and $o$ individually.
+This is framed as a ranking problem where we want the probability of the correct entity (e.g., $o$) to rank higher than all possible negative entities.
+Given the rank of the true entity among the negatives, we calculate both the Hit@K and the mean reciprocal rank (MRR), using the filtered setting following (Bordes et al., 2013).
+Following previous work (Galkin et al., 2023; Zhang and Yao, 2022; Lee et al., 2023), we evaluate against all possible negative entities.
+Baseline Methods: We consider prominent KGC methods including NBFNet (Zhu et al., 2021), RED-GNN (Zhang and Yao, 2022), NodePiece (Galkin et al., 2021), InGram (Lee et al., 2023), DEq-InGram (Gao et al., 2023), and Neural LP (Yang et al., 2017).
+We also consider the recent foundation model ULTRA (Galkin et al., 2023).
+Note that we were unable to obtain results ISDEA+, as the current code only works properly when evaluating against 50 random negatives.
+We also omit methods that have been shown to either be prohibitively expensive (e.g.,  (Teru et al., 2020; Liu et al., 2021; Mai et al., 2021)) or require the use of textual information (Gesese et al., 2022; Daza et al., 2021).
+Hyperparameters: Due to the different types and ranges of hyperparameters used for each method, the exact hyperparameters and their ranges differ by method.
+For each method, we tried to follow the recommended ranges used by the authors.
+For InGram (Lee et al., 2023) we tuned the learning rate in $\{1e^{-3},5e^{-4}\}$ and the number of entity layers in $\{2,4\}$.
+For NodePiece (Galkin et al., 2021) we tuned the learning rate from $\{1e^{-3},1e^{-4}\}$ and the loss margin from $\{15,25\}$.
+For RED-GNN (Zhang and Yao, 2022), we tuned the learning rate from $\{5e^{-3},5e^{-4}\}$ and the dropout from $\{0.1,0.3\}$.
+In their experiments, NBFNet (Zhu et al., 2021) uses the same hyperparameters across all datasets.
+We therefore do the same.
+For Neural LP (Yang et al., 2017) we also use the default set of hyperparameters that is shared across datasets.
+Training Settings: Each model was trained on either a: Tesla V100 32Gb, NVIDIA RTX A6000 48Gb, NVIDIA RTX A5000 24Gb, or Quadro RTX 8000 48Gb.
+All models were implemented in PyTorch (Paszke et al., 2019) and Torch-Geometric (Fey and Lenssen, 2019).
+5.2. Results
 Main Results: The main results for the supervised methods can be found in Tables 3 and 4.
 We observe that on nearly every dataset, NBFNet and RED-GNN are the two best models.
-This indicates that conditional MPNNs [15], the class of model in which both belong to, are necessary for strong performance in inductive KGC.
+This indicates that conditional MPNNs (Huang et al., 2024), the class of model in which both belong to, are necessary for strong performance in inductive KGC.
 Interestingly, we observe that InGram struggles in the (E, R) setting.
 This is even true when the % of new relations is high.
-This runs counter to the results on older inductive datasets [9] where InGram excelled over NBFNet and RED-GNN.
-Lastly, we observe that Neural LP sometimes fails to converge, resulting in a near zero performance and thus causing the model to have a high performance variance.
-Table 3: (E) Inductive Results (Hits@10) for supervised methods.
+This runs counter to the results on older inductive datasets (Lee et al., 2023) where InGram excelled over NBFNet and RED-GNN.
+However, this is not true for DEq-InGram, which performs consistently well under the (E, R) setting.
+We also observe that Neural LP sometimes fails to converge, resulting in a near zero performance and thus causing the model to have a high performance variance.
+Lastly, the results when evaluating via MRR can be found in Appendix A.3 and are consistent with those shown here.
+Performance Comparison of PPR vs.
+SOTA: In our original analysis, we showed that for the older inductive datasets, the performance gap between the SOTA method of each dataset and PPR is quite small.
+Specifically, the mean % difference between the two was relatively small at 25-29% (see Figure 1).
+We now perform the same analysis on the new inductive datasets.
+The results are shown in Figure 6.
+We observe that PPR generally performs much worse than the SOTA method with a mean % difference of 101%.
+Furthermore, this is more in line with the transductive datasets which have a mean % difference of 121%.
+This suggests that in addition to the decrease in raw PPR performance, the relative performance of PPR also decreases on the newer datasets.
+Table 3. (E) Inductive Results (Hits@10) for supervised methods.
 Models CoDEx-M WN18RR HetioNet
@@
 InGram 20.1 ± 3.5 38.0 ± 2.4 8.0 ± 2.9 21.9 ± 1.1 22.3 ± 2.8
+DEq-InGram 23.8 ± 1.6 62.5 ± 0.8 19.1 ± 3.1 26.5 ± 4.1 28.8 ± 3.5
 RED-GNN 35.6 ± 2.3 72.9 ± 0.4 27.7 ± 0.3 68.3 ± 3.0 85.1 ± 2.7
 NBFNet 43.6 ± 0.2 75.5 ± 0.2 29.4 ± 2.5 72.8 ± 3.8 77.2 ± 0.4
-Table 4: (E, R) Inductive Results for supervised methods. The % of new relations are in parentheses.
+Table 4. (E, R) Inductive Results for supervised methods. The % of new relations are in parentheses.
 Models FB15k-237 CoDEx-M
@@
 InGram 23.8 ± 3.0 20.2 ± 2.0 20.4 ± 3.3 15.9 ± 10.0
+DEq-InGram 35.4 ± 2.5 27.1 ± 3.5 35.2 ± 14.4 24.7 ± 0.9
 RED-GNN 21.6 ± 5.8 33.3 ± 4.2 29.2 ± 2.9 26.5 ± 10.4
 NBFNet 27.5 ± 1.8 26.2 ± 0.3 47.7 ± 11.8 17.6 ± 10.0
-Performance Comparison of PPR vs.
-SOTA: In our original analysis, we showed that for the older inductive datasets, the performance gap between the SOTA method of each dataset and PPR is quite small.
-Specifically, the mean % difference between the two was relatively small at 25-29% (see Figure 1).
-We now perform the same analysis on the new inductive datasets.
-The results are shown in Figure 6(a).
-We observe that PPR generally performs much worse than the SOTA method with a mean % difference of 99%.
-Furthermore, this is more in line with the transductive datasets which have a mean % difference of 121%.
-This suggests that in addition to the decrease in raw PPR performance, the relative performance of PPR also decreases on the newer datasets.
-Table 5: Mean performance on Old vs. New inductive datasets by Method. Red indicates a decrease in the mean performance on the newer splits.
+Figure 6. PPR vs. SOTA Hits@10 on new datasets.
+Table 5. Mean performance on Old vs. New inductive datasets by Method. Red indicates a decrease in the mean performance on the newer splits.
 Models WN18RR (E) FB15k-237 (E, R)
-Old Ind. New Ind. % Difference Old Ind. New Ind. % Difference
-PPR 66.0 46.2 -30% 21.4 10.8 -49.5%
+Old Ind. New Ind. % Diff Old Ind. New Ind. % Diff
+PPR 66.0 45.1 -31.7% 21.4 10.8 -49.5%
 Neural LP 67.6 26.4 -61.0% 16.2 20.0 +23.1%
@@
 Mean - - -40.6% - - -9.5%
-Performance Comparison on New and Old Inductive Datasets: Previously, in Table 2, we compared both the PPR performance and $\Delta\text{SPD}$ on the new and older inductive datasets.
+Performance Comparison on New and Old Inductive Datasets: In Table 2, we compared both the PPR performance and $\Delta\text{SPD}$ on the new and older inductive datasets.
 We found that both metrics tend to be much higher on the older inductive datasets, indicating that our newer splits are effective in mitigating the shortcut.
@@
 The results are shown in Table 5.
-We find that performance of most methods drops significantly across both datasets, with an average drop of 40.6% and 9.5% on WN18RR (E) and FB15k-237 (E, R), respectively.
-This suggests that removing the shortcut has a large negative effect on the performance, thereby suggesting that our new datasets are indeed harder.
-Performance of ULTRA [34]: We further compare against ULTRA, a recent foundation model designed for fully inductive KGC.
+We find that performance of most methods drops significantly on both datasets, with an average drop of 40.6% and 9.5% on WN18RR (E) and FB15k-237 (E, R), respectively.
+This suggests that removing the shortcut has a large negative effect on the performance, suggesting that our new datasets are indeed harder.
+Performance of ULTRA (Galkin et al., 2023): We further compare against ULTRA, a recent foundation model designed for fully inductive KGC.
 We evaluate ULTRA under the 0-shot setting.
-See Appendix E for more details on the versions of ULTRA used.
-The results are in Figure 6(b) where for each dataset, we average the results across the different inference graphs (full results in Appendix A.2).
+Since the setting of ULTRA differs from that of the other methods (i.e., 0-shot), we display it separately from the other methods in Tables 3 and 4.
+See Appendix D for more details on the versions of ULTRA used.
+The results are in Figure 7 where for each dataset, we average the results across the different inference graphs (full results in Appendix A.2).
 On the (E) task, ULTRA is comparable to the supervised SOTA.
 However, on the (E, R) task, ULTRA significantly outperforms other methods.
-This suggests that ULTRA contains a much greater ability than other existing methods to generalize to unseen relations.
-(a) PPR vs. SOTA Hits@10 on newer datasets
-(b) Mean Hits@10 for ULTRA vs. SOTA
-Figure 6: (a) Performance of PPR vs.
-SOTA on new inductive datasets.
-Note that we abbreviate CoDEx-M as CM, FB15k-237 as FB, HetioNet as HN and WN18RR as WN.
-Furthermore, the number represents the inference graph.
-(b) Mean performance of the foundation model ULTRA [34] vs.
-the supervised SOTA on the newer datasets.
-6 Conclusion
+This suggests that ULTRA contains a greater generalization ability than other methods.
+Figure 7. Mean Hits@10 for ULTRA vs. Supervised SOTA
+6. Conclusion
 In this paper we study the problem of constructing datasets for inductive knowledge graph completion.
-Upon examination, we find that we can achieve competitive performance on most inductive datasets through the use of Personalized PageRank [12], which ignores the relational structure of the graph.
+Upon examination, we find that we can achieve competitive performance on most inductive datasets through the use of Personalized PageRank (Page et al., 1999), which ignores the relational structure of the graph.
 Through our study, we uncover that this shortcut is due to how inductive datasets are created.
@@
 For future work, we plan to explore creating inductive KG datasets that aren’t sampled from existing transductive datasets.
+Acknowledgements.
+This research is supported by the National Science Foundation (NSF) under grant numbers CNS2321416, IIS2212032, IIS2212144, IOS2107215, DUE2234015, CNS2246050, DRL2405483 and IOS2035472, US Department of Commerce, Gates Foundation, the Michigan Department of Agriculture and Rural Development, Amazon, Meta, and SNAP.
 Appendix A Additional Results
-A.1 Supervised SOTA vs. PPR Performance
+A.1. Supervised SOTA vs. PPR Performance
 In this section we show the performance in terms of Hits@10 for the SOTA supervised method vs.
@@
 These are shown in Tables 67, and 8 for the (E), (E, R), and transductive datasets, respectively.
-For an overview of SOTA performance on additional KG datasets, please see [34].
-Table 6: Hits@10 for PPR vs. Supervised SOTA on the (E) Inductive Datasets
+For an overview of SOTA performance on additional KG datasets, please see (Galkin et al., 2023).
+Table 6. Hits@10 for PPR vs. Supervised SOTA on the (E) Inductive Datasets
 Dataset Supervised SOTA PPR % Difference SOTA Method
-WN v1 82.6 77.1 7% NBFNet [8]
-WN v2 79.8 74.4 12% NBFNet [8]
-WN v3 56.8 45.2 26% NBFNet [8]
-WN v4 74.3 67.3 10% A*Net [16]
-FB v1 60.7 41.2 47% NBFNet [8]
-FB v2 70.4 47.6 48% NBFNet [18]
-FB v3 66.7 43.5 53% NBFNet [8]
-FB v4 66.8 38.4 74% NBFNet [8]
-ILPC-S 25.1 19.8 27% NodePiece [19]
-ILPC-L 14.6 22.5 -35% NodePiece [19]
-Table 7: Hits@10 for PPR vs. Supervised SOTA on the (E, R) Inductive Datasets
+WN v1 82.6 77.1 7% NBFNet (Zhu et al., 2021)
+WN v2 79.8 74.4 12% NBFNet (Zhu et al., 2021)
+WN v3 56.8 45.2 26% NBFNet (Zhu et al., 2021)
+WN v4 74.3 67.3 10% A*Net (Zhu et al., 2023)
+FB v1 60.7 41.2 47% NBFNet (Zhu et al., 2021)
+FB v2 70.4 47.6 48% NBFNet (Shomer et al., 2023b)
+FB v3 66.7 43.5 53% NBFNet (Zhu et al., 2021)
+FB v4 66.8 38.4 74% NBFNet (Zhu et al., 2021)
+ILPC-S 25.1 19.8 27% NodePiece (Galkin et al., 2021)
+ILPC-L 14.6 22.5 -35% NodePiece (Galkin et al., 2021)
+Table 7. Hits@10 for PPR vs. Supervised SOTA on the (E, R) Inductive Datasets
 Dataset Supervised SOTA PPR % Difference SOTA Method
-FB-100 37.1 22.2 67% InGram [9]
-FB-75 32.5 21.9 48% InGram [9]
-FB-50 21.8 20.5 6% InGram [9]
-FB-25 27.1 20.9 30% InGram [9]
-WK-100 16.9 15.8 7% InGram [9]
-WK-75 36.2 29.5 23% InGram [9]
-WK-50 13.5 10.6 27% InGram [9]
-WK-25 30.9 23.2 33% InGram [9]
-Table 8: Hits@10 for PPR vs. Supervised SOTA on the Transductive Datasets
+FB-100 37.1 22.2 67% InGram (Lee et al., 2023)
+FB-75 32.5 21.9 48% InGram (Lee et al., 2023)
+FB-50 21.8 20.5 6% InGram (Lee et al., 2023)
+FB-25 27.1 20.9 30% InGram (Lee et al., 2023)
+WK-100 16.9 15.8 7% InGram (Lee et al., 2023)
+WK-75 36.2 29.5 23% InGram (Lee et al., 2023)
+WK-50 13.5 10.6 27% InGram (Lee et al., 2023)
+WK-25 30.9 23.2 33% InGram (Lee et al., 2023)
+Table 8. Hits@10 for PPR vs. Supervised SOTA on the Transductive Datasets
 Dataset Supervised SOTA PPR % Difference SOTA Method
-FB15k-237 66.6 2.7 2367% NBF+TAGNet [18]
-WN18RR 59.9 46.2 30% NBF+TAGNet [18]
-CoDEx-M 49.0 9.0 444% ComplEx RP [35]
-CoDEx-S 66.3 8.6 671% ComplEx RP [35]
-CoDEx-L 47.3 9.0 426% ComplEx RP [35]
-Hetionet 40.3 2.4 1579% RotatE [36]
-DBPedia100k 41.8 30.2 38% ComplEx-NNE+AER [37]
-A.2 Performance of ULTRA
-We include the full results of ULTRA [34] on each dataset and inference graph. The result on the (E) datasets are in Table 9 while those for the (E, R) datasets are in Table 10.
-Table 9: (E) Inductive Results for ULTRA [34]
-Metric CoDEx-M WN18RR HetioNet
-Inference 1 Inference 1 Inference 2 Inference 1 Inference 2
-MRR 30.2 64.7 21.4 57.9 72.7
-Hits@10 46.6 72.7 38.5 69.1 86.3
-Table 10: (E) Inductive Results for ULTRA [34]
+FB15k-237 66.6 2.7 2367% NBF+TAGNet (Shomer et al., 2023b)
+WN18RR 59.9 46.2 30% NBF+TAGNet (Shomer et al., 2023b)
+CoDEx-M 49.0 9.0 444% ComplEx RP (Chen et al., 2021)
+CoDEx-S 66.3 8.6 671% ComplEx RP (Chen et al., 2021)
+CoDEx-L 47.3 9.0 426% ComplEx RP (Chen et al., 2021)
+Hetionet 40.3 2.4 1579% RotatE (Sun et al., 2019)
+DBPedia100k 41.8 30.2 38% ComplEx-NNE+AER (Ding et al., 2018)
+A.2. Performance of ULTRA
+We include the full results of ULTRA (Galkin et al., 2023) on each dataset and inference graph. The result on the (E) datasets are in Table 9 while those for the (E, R) datasets are in Table 10.
+Table 9. (E) Inductive Results for ULTRA (Galkin et al., 2023)
+Dataset Graph Hits@10 MRR
+CoDEx-M Inference 1 30.2 46.6
+WN18RR Inference 1 64.7 72.7
+Inference 2 21.4 38.5
+HetioNet Inference 1 57.9 69.1
+Inference 2 72.7 86.3
+Table 10. (E) Inductive Results for ULTRA (Galkin et al., 2023)
 Metric FB15k-237 CoDEx-M
@@
 Hits@10 69.6 64.5 45.7 91.3
-A.3 Performance of Other Potential Shortcuts
-In our study, we show that PPR can achieve strong performance on most inductive KG datasets, indicating a shortcut.
-A natural question is whether this is true for just PPR, or if other shortcuts exist.
-Another bias discussed in KG literature is degree bias [38].
-In their study, they show that that KG methods tend to perform better on entities with a higher degree.
-Specifically, they observe that what matters is the degree of the entity being predicted (they refer to this as the “tail degree”).
-Based on this, we study whether the tail degree can is a good predictor of performance.
-We show that results on a number of inductive and transductive datasets in Table 11.
-We report Hits@10 for all results.
-We find that the tail degree often achieves a much lower performance than PPR.
-Specifically, for the datasets in Table  11, while the PPR performance is on average only 31% lower than the SOTA, the tail degree performance is 84% lower.
-This indicates the severity of the PPR shortcut is much more severe as compared to other known biases like the tail degree.
-Table 11: Performance of Tail Degree [38], PPR, and SOTA method. Performance is in terms of Hits@10. We further bold the larger of the PPR and tail degree performance.
+Table 11. Performance of Tail Degree (Shomer et al., 2023a), PPR, and SOTA method. Performance is in terms of Hits@10. We further bold the larger of the PPR and tail degree performance.
 Task Dataset Tail Degree PPR SOTA
@@
 DBPedia100k 2.9 30.2 41.8
+A.3. MRR Performance
+We further include the performance on the new datasets when evaluating via MRR.
+The results are shown in Tables 12 and 13 for the (E) and (E, R) results, respectively.
+We can observe that these results are highly consistent when evaluating via Hits@10.
+That is, under the (E) setting both NBFNet and RED-GNN perform best, while under the (E, R) setting DEq-InGram also excels.
+This suggests that our observations hold under multiple evaluation metrics.
+Table 12. (E) Inductive Results (MRR) for supervised methods.
+Models CoDEx-M WN18RR HetioNet
+Inference 1 Inference 1 Inference 2 Inference 1 Inference 2
+PPR 3.3 25.6 7.9 2.3 2.1
+Neural LP 9.2 ± 11.2 23.9 ± 2.0 10.5 ± 1.8 7.3 ± 9.9 6.7 ± 9.3
+NodePiece 3.1 ± 0.3 20.6 ± 0.8 3.3 ± 0.4 5.1 ± 0.6 7.2 ± 0.6
+InGram 10.1 ± 1.5 15.9 ± 0.7 3.8 ± 1.1 11.2 ± 0.5 11.5 ± 2.1
+DEq-InGram 11.2 ± 0.6 25.4 ± 0.5 9.8 ± 1.7 16.5 ± 2.9 16.9 ± 2.0
+RED-GNN 22.1 ± 1.8 67.8 ± 0.5 16.4 ± 1.3 53.8 ± 2.5 68.6 ± 4.3
+NBFNet 27.6 ± 0.2 66.6 ± 0.1 17.6 ± 0.1 39.1 ± 1.2 45.2 ± 0.5
+Table 13. (E, R) Inductive Results (MRR) for supervised methods. The % of new relations are in parentheses.
+Models FB15k-237 CoDEx-M
+Inference 1 (27%) Inference 2 (63%) Inference 1 (10%) Inference 2 (57%)
+PPR 3.8 4.7 3.3 7.1
+Neural LP 8.8 ± 4.9 11.3 ± 6.6 13.8 ± 18.8 7.4 ± 9.4
+NodePiece 1.4 ± 0.1 2.6 ± 0.2 1.5 ± 0.2 1.6 ± 0.4
+InGram 11.2 ± 1.9 9.2 ± 0.9 10.3 ± 1.2 7.2 ± 4.2
+DEq-InGram 19.5 ± 2.0 13.4 ± 1.2 18.7 ± 7.6 13.0 ± 1.3
+RED-GNN 11.8 ± 2.8 19.8 ± 2.2 16.8 ± 1.6 14.7 ± 5.3
+NBFNet 15.8 ± 0.9 15.9 ± 0.2 27.9 ± 8.8 8.5 ± 4.8
+A.4. Performance of Other Potential Shortcuts
+In our study, we show that PPR can achieve strong performance on most inductive KG datasets, indicating a shortcut.
+A natural question is whether this is true for just PPR, or if other shortcuts exist.
+Another bias discussed in KG literature is degree bias (Shomer et al., 2023a).
+In their study, they show that that KG methods tend to perform better on entities with a higher degree.
+Specifically, they observe that what matters is the degree of the entity being predicted (they refer to this as the “tail degree”).
+Based on this, we study whether the tail degree can is a good predictor of performance.
+We show that results on a number of inductive and transductive datasets in Table 11.
+We report Hits@10 for all results.
+We find that the tail degree often achieves a much lower performance than PPR.
+Specifically, for the datasets in Table  11, while the PPR performance is on average only 31% lower than the SOTA, the tail degree performance is 84% lower.
+This indicates the severity of the PPR shortcut is much more severe as compared to other known biases like the tail degree.
 Appendix B Datasets
-B.1 Existing Datasets
-We detail the statistics of all existing transductive and inductive datasets in Tables 12, 13, and 14, respectively.
-We further include the licenses for each in Table 15.
-Note that we omit YAGO3-10 [39] as [40] show that the dataset is dominated by two duplicate relations, making most triples trivial to classify.
-Also, we omit NELL-995 and any inductive datasets derived from it due to findings by [31] that show that most triples in the dataset are either too generic or meaningless.
-Table 12: Statistics for Transductive Datasets.
+B.1. Existing Datasets
+We detail the statistics of all existing transductive and inductive datasets in Tables 14, 15, and 16, respectively.
+Note that we omit YAGO3-10 (Mahdisoltani et al., 2013) as (Akrami et al., 2020) show that the dataset is dominated by two duplicate relations, making most triples trivial to classify.
+Also, we omit NELL-995 and any inductive datasets derived from it due to findings by (Safavi and Koutra, 2020) that show that most triples in the dataset are either too generic or meaningless.
+Table 14. Statistics for Transductive Datasets.
 Dataset #Entities #Relations #Train #Validation #Test
-FB15k-237 [33] 14,541 237 272,115 17,535 20,466
-WN18RR [27] 40,943 11 86,835 3,034 3,134
-CoDEx-S [31] 2,034 42 32,888 1,827 1,828
-CoDEx-M [31] 17,050 51 185,584 10,310 10,311
-CoDEx-L [31] 77,951 69 551,193 30,622 30,622
-HetioNet [32] 45,158 24 2,025,177 112,510 112,510
-DBpedia100k [37] 99,604 470 597,572 50,000 50,000
-Table 13: Statistics for (E) Inductive Datasets.
+FB15k-237 (Toutanova and Chen, 2015) 14,541 237 272,115 17,535 20,466
+WN18RR (Dettmers et al., 2018) 40,943 11 86,835 3,034 3,134
+CoDEx-S (Safavi and Koutra, 2020) 2,034 42 32,888 1,827 1,828
+CoDEx-M (Safavi and Koutra, 2020) 17,050 51 185,584 10,310 10,311
+CoDEx-L (Safavi and Koutra, 2020) 77,951 69 551,193 30,622 30,622
+HetioNet (Himmelstein et al., 2017) 45,158 24 2,025,177 112,510 112,510
+DBpedia100k (Ding et al., 2018) 99,604 470 597,572 50,000 50,000
+Table 15. Statistics for (E) Inductive Datasets.
 Dataset #Relations Train Graph Validation Graph Test Graph
 #Entities #Triples #Entities #Triples #Valid #Entities #Triples #Test
-FB15k-237 v1 [7] 180 1,594 4,245 1,594 4,245 489 1,093 1,993 205
-FB15k-237 v2 [7] 200 2,608 9,739 2,608 9,739 1,166 1,660 4,145 478
-FB15k-237 v3 [7] 215 3,668 17,986 3,668 17,986 2,194 2,501 7,406 865
-FB15k-237 v4 [7] 219 4,707 27,203 4,707 27,203 3,352 3,051 11,714 1,424
-WN18RR v1 [7] 9 2,746 5,410 2,746 5,410 630 922 1,618 188
-WN18RR v2 [7] 10 6,954 15,262 6,954 15,262 1,838 2,757 4,011 441
-WN18RR v3 [7] 11 12,078 25,901 12,078 25,901 3,097 5,084 6,327 605
-WN18RR v4 [7] 9 3,861 7,940 3,861 7,940 934 7,084 12,334 1,429
-ILPC-S [10] 48 10,230 78,616 6,653 20,960 2,908 6,653 20,960 2902
-ILPC-L [10] 65 46,626 202,446 29,246 77,044 10,179 29,246 77,044 10,184
-Table 14: Statistics for (E, R) Inductive Datasets.
+FB15k-237 v1 (Teru et al., 2020) 180 1,594 4,245 1,594 4,245 489 1,093 1,993 205
+FB15k-237 v2 (Teru et al., 2020) 200 2,608 9,739 2,608 9,739 1,166 1,660 4,145 478
+FB15k-237 v3 (Teru et al., 2020) 215 3,668 17,986 3,668 17,986 2,194 2,501 7,406 865
+FB15k-237 v4 (Teru et al., 2020) 219 4,707 27,203 4,707 27,203 3,352 3,051 11,714 1,424
+WN18RR v1 (Teru et al., 2020) 9 2,746 5,410 2,746 5,410 630 922 1,618 188
+WN18RR v2 (Teru et al., 2020) 10 6,954 15,262 6,954 15,262 1,838 2,757 4,011 441
+WN18RR v3 (Teru et al., 2020) 11 12,078 25,901 12,078 25,901 3,097 5,084 6,327 605
+WN18RR v4 (Teru et al., 2020) 9 3,861 7,940 3,861 7,940 934 7,084 12,334 1,429
+ILPC-S (Galkin et al., 2022) 48 10,230 78,616 6,653 20,960 2,908 6,653 20,960 2902
+ILPC-L (Galkin et al., 2022) 65 46,626 202,446 29,246 77,044 10,179 29,246 77,044 10,184
+Table 16. Statistics for (E, R) Inductive Datasets.
 Dataset Train Graph Validation Graph Test Graph
 #Entities #Rels #Triples #Entities #Rels #Triples #Valid #Entities #Rels #Triples #Test
-FB-25 [9] 5,190 163 91,571 4,097 216 17,147 5,716 4,097 216 17,147 5,716
-FB-50 [9] 5,190 153 85,375 4,445 205 11,636 3,879 4,445 205 11,636 3,879
-FB-75 [9] 4,659 134 62,809 2,792 186 9,316 3,106 2,792 186 9,316 3,106
-FB-100 [9] 4,659 134 62,809 2,624 77 6,987 2,329 2,624 77 6,987 2,329
-WK-25 [9] 12,659 47 41,873 3,228 74 3,391 1,130 3,228 74 3,391 1,131
-WK-50 [9] 12,022 72 82,481 9,328 93 9,672 3,224 9,328 93 9,672 3,225
-WK-75 [9] 6,853 52 28,741 2,722 65 3,430 1,143 2,722 65 3,430 1,144
-WK-100 [9] 9,784 67 49,875 12,136 37 13,487 4,496 12,136 37 13,487 4,496
-Table 15: Dataset Licenses.
-Datasets License
-FB15k-237 CC-BY-4.0
-WN18RR MIT License
-CoDEx-S/M/L MIT License
-HetioNet CC0 1.0 Universal
-DBpedia100k None
-FB15k-237 v1/v2/v3/v4 None
-WN18RR v1/v2/v3/v4 None
-ILPC-S/L MIT License
-FB-25/50/75/100 None
-WK-25/50/75/100 None
-B.2 New Datasets
-We further detail the statistics of all the new datasets in Tables 16 and 17.
-Table 16: Statistics for New (E) Inductive Datasets.
+FB-25 (Lee et al., 2023) 5,190 163 91,571 4,097 216 17,147 5,716 4,097 216 17,147 5,716
+FB-50 (Lee et al., 2023) 5,190 153 85,375 4,445 205 11,636 3,879 4,445 205 11,636 3,879
+FB-75 (Lee et al., 2023) 4,659 134 62,809 2,792 186 9,316 3,106 2,792 186 9,316 3,106
+FB-100 (Lee et al., 2023) 4,659 134 62,809 2,624 77 6,987 2,329 2,624 77 6,987 2,329
+WK-25 (Lee et al., 2023) 12,659 47 41,873 3,228 74 3,391 1,130 3,228 74 3,391 1,131
+WK-50 (Lee et al., 2023) 12,022 72 82,481 9,328 93 9,672 3,224 9,328 93 9,672 3,225
+WK-75 (Lee et al., 2023) 6,853 52 28,741 2,722 65 3,430 1,143 2,722 65 3,430 1,144
+WK-100 (Lee et al., 2023) 9,784 67 49,875 12,136 37 13,487 4,496 12,136 37 13,487 4,496
+B.2. New Datasets
+We further detail the statistics of all the new datasets in Tables 17 and 18.
+Table 17. Statistics for New (E) Inductive Datasets.
 Dataset Graph # Edges # Entities # Rels # Valid/Test $\Delta\text{SPD}$
@@
 Inference 1 18,258 8,660 10 1,831 4.87
-Inference 2 5,838 2975 8 572 1.47
+Inference 2 5,838 2,975 8 572 1.47
 HetioNet Train 101,667 3,971 14 11,271 NA
 Inference 1 49,590 2,279 11 5,490 0.38
-Inference 2 37,927 2455 12 4,187 0.19
-Table 17: Statistics for New (E, R) Inductive Datasets.
+Inference 2 37,927 2,455 12 4,187 0.19
+Table 18. Statistics for New (E, R) Inductive Datasets.
 Dataset Graph # Edges # Entities # Rels # Valid/Test % New Rels $\Delta\text{SPD}$
-FB15k-237 Train 45,597 2,869 105 5,062 NA NA
-Inference 1 35,937 1,835 143 3,992 62.8% 0.63
+FB15k-237 Train 45,597 2,864 104 5,062 NA NA
+Inference 1 35,937 1,835 72 3,898 62.8% 0.63
 Inference 2 51,693 2,606 143 5,735 27.1% 0.34
@@
 Inference 1 70,137 7,938 39 7,794 9.9% 0.24
-Inference 2 8,821 2,606 28 979 56.8% 0.60
-Appendix C Inductive Generation Experiments
-We give more details on the experiments conducted in Section 3.3 and shown in Figures 4.
-Given the transductive dataset FB15k-237 [33], we generate a number of different inductive datasets using the common procedure used by [7, 9].
-We describe this procedure in detail in Section 2.
-Note that for simplicity, we only created datasets for the (E) inductive task.
-However, the same conclusions should hold for (E, R).
-Lastly, we only create one graph for inference as is done in previous work.
-We generate the inductive datasets by modifying the following set of parameters:
-  • # of Initial Train Entities: This is the number of entities that are first assigned to the train graph.
-  • # of Initial Inference Entities: This is the number of entities that are first assigned to the inference graph.
-  • Max Train Neighborhood Size: We extract the 2-hop neighborhood for each of the initial train entities.
-To avoid exponential growth, we limit the number of entities selected in each hop to 50.
-For example, if node 1 has 30 1-hop neighbors and 120 2-hop neighbors, then $30+50=80$ entities are added through the expansion of node 1.
-  • Max Inference Neighborhood Size: This follows the same logic as for train but for inference.
-The default neighborhood size is set to 50 for the inductive datasets created by [7] and [9].
-For [9] they consider 10 and 20, initial entities for train and inference, respectively.
-This information is not reported for [7].
-To simulate the effect of the different parameter values, we create inductive datasets by using a number of different combinations of parameters.
-For each parameter, we modify it while holding the others constant.
-This allows us to explore the effect of just that variable.
-It allows us to avoid running an excessive amount of experiments.
-The default values are those used in [9] and noted above.
-In total, there are 17 parameter configurations.
-For each parameter configuration, we create 3 different datasets through the use of 3 different random seeds.
-This results in a total of 51 inductive datasets.
-We then evaluate the performance of PPR on each inference graph and calculate the $\Delta\text{SPD}$.
-The results for those datasets with the same configuration are then averaged together.
-All possible configurations and their resulting mean statistics can be found in Table 18.
-These results are visualized in the main content in Figures 4.
-Table 18: FB15k-237 - Inductive Generative Experiments Configurations. Results are over 3 Random Seeds.
+Inference 2 8,821 474 28 979 56.8% 0.60
+Table 19. FB15k-237 - Inductive Generative Experiments Configurations. Results are over 3 Random Seeds.
 # Train Ents # Inf. Ents Max Train Max Inf. # Train Edges # Test Edges $\Delta\text{SPD}$ PPR Hits@10
@@
 10 160 50 50 91,843 55,233 1.93 0.259
-Appendix D Additional Experimental Settings
-Training Settings: Each model was trained on either a: Tesla V100 32Gb, NVIDIA RTX A6000 48Gb, NVIDIA RTX A5000 24Gb, or Quadro RTX 8000 48Gb.
-All models were implemented in PyTorch [41] and Torch-Geometric [42].
-Hyperparameters: Due to the different types and ranges of hyperparameters used for each method, the exact hyperparameters and their ranges differ by method.
-For each method, we tried to follow the recommended ranges used by the authors.
-For InGram [9] we tuned the learning rate in $\{1e^{-3},5e^{-4}\}$ and the number of entity layers in $\{2,4\}$.
-For NodePiece [19] we tuned the learning rate from $\{1e^{-3},1e^{-4}\}$ and the loss margin from $\{15,25\}$.
-For RED-GNN [11], we tuned the learning rate from $\{5e^{-3},5e^{-4}\}$ and the dropout from $\{0.1,0.3\}$.
-In their experiments, NBFNet [8] uses the same hyperparameters across all datasets.
-We therefore do the same.
-For Neural LP [13] we also use the default set of hyperparameters that is shared across datasets.
-Evaluation: For a given test triple $(s,r,o)$, we strive to predict both $s$ and $o$ individually.
-This is framed as a ranking problem where we want the probability of the correct entity (e.g., $o$) to rank higher than all possible negative entities.
-We can calculate the rank of the true entity over all negatives by the following where $p(\cdot)$ denotes the model probability for a triple:
-[…]
-We note that a lower rank, indicates better performance.
-In practice, some of the negative entities will be true triples observed during training.
-As such, following [4] we use the filtered setting and remove those entities from the ranking.
-Once we obtain the rank of the true entity (i.e., “how many negative entities does it have a higher probability than”) we compute the Hits@K%.
-Hits@Kcomputes the percentage of sample where the following inequality holds true, $\text{Rank}\leq K$.
-Appendix E ULTRA Settings
-When evaluating ULTRA [34] we use the 0-shot setting.
-By default, we use the checkpoint they provided that was trained on the transductive datasets: FB15k-237 [33], WN18RR [27], CoDEx-M [31], and NELL-995 [43].
+Appendix C Inductive Generation Experiments
+We give more details on the experiments conducted in Section 3.3 and shown in Figures 4.
+Given the transductive dataset FB15k-237 (Toutanova and Chen, 2015), we generate a number of different inductive datasets using the common procedure used by (Teru et al., 2020; Lee et al., 2023).
+We describe this procedure in detail in Section 2.
+Note that for simplicity, we only created datasets for the (E) inductive task.
+However, the same conclusions should hold for (E, R).
+Lastly, we only create one graph for inference as is done in previous work.
+We generate the inductive datasets by modifying the following set of parameters:
+  • # of Initial Train Entities: This is the number of entities that are first assigned to the train graph.
+  • # of Initial Inference Entities: This is the number of entities that are first assigned to the inference graph.
+  • Max Train Neighborhood Size: We extract the 2-hop neighborhood for each of the initial train entities.
+To avoid exponential growth, we limit the number of entities selected in each hop to 50.
+For example, if node 1 has 30 1-hop neighbors and 120 2-hop neighbors, then $30+50=80$ entities are added through the expansion of node 1.
+  • Max Inference Neighborhood Size: This follows the same logic as for train but for inference.
+The default neighborhood size is set to 50 for the inductive datasets created by (Teru et al., 2020) and (Lee et al., 2023).
+For (Lee et al., 2023) they consider 10 and 20, initial entities for train and inference, respectively.
+This information is not reported for (Teru et al., 2020).
+To simulate the effect of the different parameter values, we create inductive datasets by using a number of different combinations of parameters.
+For each parameter, we modify it while holding the others constant.
+This allows us to explore the effect of just that variable.
+It allows us to avoid running an excessive amount of experiments.
+The default values are those used in (Lee et al., 2023) and noted above.
+In total, there are 17 parameter configurations.
+For each parameter configuration, we create 3 different datasets through the use of 3 different random seeds.
+This results in a total of 51 inductive datasets.
+We then evaluate the performance of PPR on each inference graph and calculate the $\Delta\text{SPD}$.
+The results for those datasets with the same configuration are then averaged together.
+All possible configurations and their resulting mean statistics can be found in Table 19.
+These results are visualized in the main content in Figures 4.
+Appendix D ULTRA Settings
+When evaluating ULTRA (Galkin et al., 2023) we use the 0-shot setting.
+By default, we use the checkpoint they provided that was trained on the transductive datasets: FB15k-237 (Toutanova and Chen, 2015), WN18RR (Dettmers et al., 2018), CoDEx-M (Safavi and Koutra, 2020), and NELL-995 (Xiong et al., 2017).
 However, since we create our own inductive splits from FB15k-237, CoDEx-M, and WN18RR, there is risk of test leakage.
@@
 We follow the same settings used to train the original checkpoints provided by the authors of ULTRA. Lastly, since HetioNet is not one of the four datasets used, no additional model is needed.
-The results are shown in Table 19, where the “W/o” column represents the results when removing the parent transductive dataset and “With” corresponds to the original pre-trained checkpoint provided by the authors of ULTRA [34].
+The results are shown in Table 22, where the “W/o” column represents the results when removing the parent transductive dataset and “With” corresponds to the original pre-trained checkpoint provided by the authors of ULTRA (Galkin et al., 2023).
 In practice, we observe a small but noteworthy decrease in performance when removing the offending transductive dataset from the pre-trained model.
-Table 19: Performance (Hits@10) of ULTRA when pre-trained and w/o the parent transductive dataset. Note that this doesn’t apply to HetioNet.
+Table 20. PPR Performance on PediaTypes (Gao et al., 2023) datasets.
+Dataset 50 Negatives All Negatives
+Neural Hits@10 PPR Hits@10 % Difference PPR Hits@10 $\Delta\text{SPD}$
+EN-FR 95.4% 96.0% -0.6% 44.1% 2.7
+FR-EN 90.6% 96.2% -5.8% 48.1% 2.3
+EN-DE 97.7% 93.9% 4.0% 34.7% 2.1
+DE-EN 95.0% 90.6% 4.9% 20.8% 1.6
+DB-WD 86.6% 76.4% 13.4% 23.5% 0.6
+WD-DB 91.5% 92.0% -0.5% 42.9% 2.5
+DB-YG 71.5% 76.9% -7.0% 15.2% 2.7
+YG-DB 80.5% 85.1% -5.4% 30.7% 3.5
+Table 21. PPR Performance on WikiTopics (Gao et al., 2023) datasets.
+Dataset 50 Negatives All Negatives
+Neural Hits@10 PPR Hits@10 % Difference PPR Hits@10 $\Delta\text{SPD}$
+Art 92.3% 79.3% 16.4% 14.7% 1.1
+Award 93.6% 95.8% -2.3% 24.2% 3.4
+Edu 97.0% 72.4% 34.0% 20.8% 2.7
+Health 98.9% 91.5% 8.1% 59.6% 3.3
+Infra 97.8% 97.2% 0.6% 61.4% 5.4
+Sci 96.7% 84.3% 14.7% 25.5% 4.4
+Sport 81.2% 82.9% -2.1% 14.4% 0.3
+Tax 93.7% 64.1% 46.1% 29.0% 0.8
+Loc N/A 96.5% N/A 74.1% 1.4
+Org N/A 94.0% N/A 8.9% 2.8
+People N/A 94.9% N/A 64.6% 2.7
+Table 22. Performance (Hits@10) of ULTRA when pre-trained and w/o the parent transductive dataset. Note that this doesn’t apply to HetioNet.
 Dataset Inference Graph W/o With % Difference
@@
 2 91.1 91.3 -0.2%
-Appendix F Implementation
-All code used for experiments in this study can be found in the following anonymized repository – https://github.com/HarryShomer/Better-Inductive-KGC.
-Please see the README in the repository for more information on how to run the different experiments.
-Appendix G Limitations
-One potential limitation of our study is that the new inductive datasets are still being created from existing transductive datasets.
-While our method can help create better and more realistic inductive scenarios, this is limiting as we are still reliant on good quality transductive datasets.
-Future work can focus on creating inductive KGC datasets directly from existing large-scale knowledge graphs, thereby bypassing the need to sample from existing transductive datasets.
-Appendix H Impact Statement
-Our method contributed positively to the field of knowledge graph reasoning.
-By introducing newer and better inductive datasets, we better align the task of KGC to it’s real-world applications.
-This is essential, because if we want to perform inductive KGC in real-world tasks, we have to be certain that the methods actually work well.
-As such, the new datasets can serve as a stronger barometer of actual performance on inductive KGC and will help spur the development of newer and more effective KGC techniques.
-Since KGC has applications in many different fields including question answering, biology, and recommender systems, there is a strong need for methods that can perform well.
-We have also carefully considered the broader impact of our work from different perspectives and find that there is no apparent risk.
+Appendix E Analysis of the WikiTopics and PediaTypes Datasets
+We further include an analysis of the WikiTopics and PediaTypes inductive datasets introduced by Gao et al.
+(2023).
+For these datasets, the train and inference graphs contain completely separate entities and relations.
+For each dataset we record the (a) Neural SOTA performance (taken from (Gao et al., 2023), (b) PPR performance, (c) % difference in performance, and (d) $\Delta\text{SPD}$.
+In their original study, Gao et al.
+(2023) use 50 random negatives during evaluation, instead of all negative entities used in ours.
+To account for this, we compute the PPR performance using both evaluation settings.
+The results for the PediaTypes and WikiTopics datasets are in Tables 20 and 21, respectively.
+From the results we can make multiple observations.
+First, PPR performs well under both settings.
+Under the all negatives setting, the mean PPR performance is 32.5% and 38.3% on PediaTypes and WikiTopics, respectively.
+Second, under 50 negatives, there is little difference in the PPR and Neural performance.
+Specifically, the % difference in performance is 0.4% and 14.2%, respectively, on PediaTypes and WikiTopics.
+This suggests that neural methods can only slightly outperform PPR under this setting.
+Lastly, the $\Delta\text{SPD}$ is high on both sets of datasets.
+This further corroborates our findings in Section 3.
+This analysis shows that the WikiTopics and PediaTopics datasets introduced by Gao et al. (2023) fall prey to the same shortcut found in other inductive datasets studied in Section 3.